### v1

# One-Shot Graph Representation Learning Using Hyperdimensional Computing

Abstract
We present a novel, simple, fast, and efficient approach for semi-supervised learning on graphs. The proposed approach takes advantage of hyper-dimensional computing which encodes data samples using random projections into a high dimensional space (HD space for short). Specifically, we propose a Hyper-dimensional Graph Learning (HDGL) algorithm that leverages the injectivity property of the node representations of a family of graph neural networks. HDGL maps node features to the HD space and then uses HD operators such as bundling and binding to aggregate information from the local neighborhood of each node. Results of experiments with widely used benchmark data sets show that HDGL achieves predictive performance that is competitive with the state-of-the-art deep learning methods, without the need for computationally expensive training.
1 Introduction
Social networks, citation networks, etc., are naturally represented as graphs where nodes represent the entities that make up the network and edges between pairs of nodes encode relationships between entities. For example, in a social network, nodes represent individuals and links denote social connections; in a citation network, nodes represent articles, and (directed) links denote citations. Over the past decade, there has been considerable interest in graph learning problems such as node classification and link prediction Li and Pi (2020); Kumar et al. (2020). Such problems are typically solved using graph embedding techniques Zhou et al. (2022); Cai et al. (2018); Xu (2021); Wang et al. (2022).
Graph Neural Networks (GNNs), originally introduced by Sperduti and Starita (1997); Frasconi et al. (1998); Gori et al. (2005); Scarselli et al. (2009), which learn low-dimensional embeddings of nodes in a graph by aggregrating information from local neighborhoods of the nodes, have recently emerged as the dominant approach to graph representation learning Kipf and Welling (2017); Veličković et al. (2018); Hamilton et al. (2017), because of their superior performance across a broad spectrum of applications Wu et al. (2020); Zhang et al. (2018); Zhou et al. (2022).
Like other deep learning models, GNNs are data hungry, require extensive hyperparameter tuning, and achieve their superior performance at considerable computational cost, resulting in a large carbon footprint with attendant environmental impact Desislavov et al. (2023); Strubell et al. (2020); de Vries (2023); Thompson et al. (2021). Hence, there is an urgent need for computationally more efficient, and consequently, environmentally more sustainable alternatives to deep learning models in general, with a specific emphasis on graph neural networks.
Hyperdimensional computing Kanerva (1988, 2009); Neubert et al. (2019); Thomas et al. (2021) offers a brain-inspired alternative to deep neural networks for artificial intelligence and machine learning applications, that maps each object using a sufficiently high-dimensional random encoding (HD encoding for short) to produce a binary or bipolar vector Kanerva (1988, 2009). All computations are performed on the objects thus encoded using simple operations, e.g., element-wise additions and dot products Kanerva (1988); Neubert et al. (2019); Thomas et al. (2021). Computations on HD vectors, because they yield binary or bipolar vectors, can be realized using low-precision, fast, low-power, energy-efficient hardware Kleyko et al. (2023a, b). Hence, there is a growing interest in the use of HD computing in artificial intelligence and machine learning Neubert et al. (2019); Kleyko et al. (2023b); Ge and Parhi (2020); Morris et al. (2022).
Against this background, we investigate Hyper-dimensional Graph Learner (HDGL), a one-shot approach to graph learning, that makes use of hyper-dimensional computing Kanerva (1988, 2009); Neubert et al. (2019); Thomas et al. (2021). Specifically, we outline a methodology for designing an HD mapping for node embedding that achieves competitive performance in semi-supervised node labeling tasks when compared to GNNs, which necessitate considerably more costly training. The key contributions of the paper are as follows:
  • We show that how to design a HD node embeddings that match the expressive power of GNN embeddings.
  • We propose HDGL, a graph learner that takes advantage of carefully designed hyperdimensional node embeddings to perform one-shot learning on graphs for semi-supervised node labeling tasks.
  • We present results of extensive experimental comparisons of HDGL with the state-of-the-art GNN methods on several standard benchmark data sets which show that HDGL achieves performance that is competitive with that of GNN, while avoiding the need for computationally expensive training.
The rest of the paper is organized as follows: Section 2 reviews the HD computing framework, its basic operations, and their properties, and reviews the expressive power of GNN. Section 3 introduces HDGL, the proposed HD computing based one-shot algorithm for learning node embeddings. Section 4 describes results of extensive experiments that compare HDGL with the state-of-the-art GNN-based node embedding methods. Section 5 concludes with a brief summary of the paper and an outline of some directions for further research.
2 Background and Preliminaries
2.1 Node Embedding Problem
Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ represents set of undirected nodes, $\mathcal{E}$ represents the set of edges $\mathcal{E}\subseteq\{(u,v)|u,v,\in\mathcal{V}\}$. Let $N=|\mathcal{V}|$ denote the number of nodes in the graph. Supposed each node $v\in\mathcal{V}$ is described by a tuple of $d$ features $\mathbf{x}_{v}\in\mathbb{R}^{d}$. Let $y_{v}\in\{1,\cdots,L\}$ be the the label associated with each $v\in\mathcal{V}$. The $k$-hop neighbors of node $v$ in graph $\mathcal{G}$ are represented using multi-set $\mathcal{N}^{k}(v)$. The goal is to learn a latent representation $\phi(v)$ of each node $v$ using the structure of $\mathcal{G}$ and $\mathbf{x}_{v}$ to accurately predict $y_{v}$.
2.2 Properties of HD Computing
Hyper-dimensional computing, originally introduced by Kanerva (1988) is a brain-inspired approach to cognitive computing that encodes each object $x\in\mathcal{X}$ using a high dimensional static mapping $\phi:\mathcal{X}\rightarrow\mathcal{H}$, where $\mathcal{H}$ is typically $\{-1,1\}^{n}$ or $\{0,1\}^{n}$ where $n$ is sufficiently large Kanerva (1988, 2009). All computations are performed in $\mathcal{H}$, using simple operations, e.g., element-wise additions and dot products Kanerva (1988); Neubert et al. (2019). The mapping $\phi$ is often random, which because its elements are low precision, lends itself to fast, low-power, energy-efficient hardware realizations Kleyko et al. (2023a, b). In this paper, we assume that =$\mathcal{H}=\{0,1\}^{n}$. We proceed to briefly summarize some of the key properties of HD representation and operations that can be used to compute with it.
Near Orthogonality of Random vectors in HD Space. HD encodings of objects are produced by sampling each dimension independently and uniformly from a distribution. Let $\mathbf{a}\in\{0,1\}^{n}$ and $\mathbf{b}\in\{0,1\}^{n}$ be two such HD vectors where $a_{i}\sim Ber(p=0.5);\forall i\in\{1,\cdots,n\}$ and $b_{i}\sim Ber(p=0.5);\forall i\in\{1,\cdots,n\}$. We can then say that $\mathbf{a}$ and $\mathbf{b}$ are near orthogonal or $d_{H}[\mathbf{a},\mathbf{b}]$, the normalized hamming distance between $\mathbf{a}$ and $\mathbf{b}$, $\approx 0.5$ Kanerva (1988). This is also known as a blessing of HD spaces.
High Memory Capacity of HD Representation. As noted by Neubert et al. (2019), in the case of an $n$-dimensional binary HD space where each of the $n$-dimensional vectors is $f$-sparse, e.g., the number of nonzero entries is $fn$, the resulting HD space has capacity given by $\binom{n}{\lfloor f\cdot n\rfloor}$. For example, when $n=1000$, and $f=0.05$, the resulting HD space can store approximately $10^{80}$ or almost as many patterns as the number of atoms in the observable universe (which is estimated to be between $10^{78}$ and $10^{82}$).
Noise or Error Tolerance. Due to the distributed representation of data across numerous dimensions in HD spaces, errors in HD computing are limited to only a fraction of bits Ahmad and Hawkins (2015); Neubert et al. (2019); Thomas et al. (2021). Thus, HD computing offers noise tolerance without the need for expensive error correction mechanisms Neubert et al. (2019).
2.3 HD Computing Operations
Bundling, denoted by $\oplus$, is a bitwise majority operation represented by :$\oplus:\{0,1\}^{n}\times\{0,1\}^{n}\rightarrow\{0,1\}^{n}$. Bundling two randomly sampled HD vectors results in a new vector which is similar to the input vectors. Bundling $\oplus$ is associative and commutative i.e $\mathbf{a}\oplus(\mathbf{b}\oplus\mathbf{c})=(\mathbf{a}\oplus\mathbf{b})\oplus\mathbf{c}=(\mathbf{c}\oplus\mathbf{a})\oplus\mathbf{b}$. Note that in case of Bundling an even number of HD vectors, there is a possibility of ties since Bundling is a bitwise majority operation. To break ties, a tie-breaking policy, e.g., random choice of 0 or 1 on dimensions where ties occur. Bundling can be thought of as the HD space analog of the mean operation euclidean spaces. Bundling, because it commutes, can be used to represent sets as well as multisets of objects encoded using HD vectors.
Binding, denoted by $\otimes$, is a bitwise Exclusive OR operation represented by :$\otimes:\{0,1\}^{n}\times\{0,1\}^{n}\rightarrow\{0,1\}^{n}$, is an operation which takes two input vectors and maps them to a new vector. An important property of Binding is that Binding two HD vectors results in a vector that is dissimilar to both. Formally, let $\mathbf{a}\in\{0,1\}^{n}$ and $\mathbf{b}\in\{0,1\}^{n}$ be two sampled HD vectors and $\mathbf{c}=\mathbf{a}\otimes\mathbf{b}$. Due to the near-orthogonality of HD vectors, we can then say that $d_{H}[\mathbf{c},\mathbf{a}]\approx 0.5$ and $d_{H}[\mathbf{c},\mathbf{b}]\approx 0.5$. Binding is invertible, associative and commutative i.e $\mathbf{b}\otimes(\mathbf{a}\otimes\mathbf{b})=(\mathbf{b}\otimes\mathbf{a})\otimes\mathbf{b}=\mathbf{a}\otimes\mathbf{b}\otimes\mathbf{b}=\mathbf{a}$. Moreover, Binding is distributive over bundling i.e $\mathbf{a}\otimes(\mathbf{b}\oplus\mathbf{c})=(\mathbf{a}\otimes\mathbf{b})\oplus(\mathbf{a}\otimes\mathbf{c})$. Note that Binding also has a reflective property i.e the distance between two vectors doesn’t change when both vectors are bound to the same vector. Formally, $d_{H}[\mathbf{a},\mathbf{b}]=d_{H}[\mathbf{c}\otimes\mathbf{a},\mathbf{c}\otimes\mathbf{b}]$. Note that this also works even vectors are slightly different i.e $d_{H}[\mathbf{a},\mathbf{b}]\approx d_{H}[\mathbf{c}\otimes\mathbf{a}^{\prime},\mathbf{c}\otimes\mathbf{b}]$ if $\mathbf{a}\approx\mathbf{a}^{\prime}$.
Rotation, also called permutation, denoted by $\Pi$, is a bitwise left shift operation $\Pi:\{0,1\}^{n}\rightarrow\{0,1\}^{n}$. A key property of Rotation is that rotation results in dissimilar/orthogonal vectors i.e $d_{H}[\mathbf{a},\Pi(\mathbf{a})]\approx 0.5$ and it is distributive over both Bundling and Binding and like binding, rotation is also reflective i.e $d_{H}[\mathbf{a},\mathbf{b}]\approx d_{H}[\Pi(\mathbf{a}),\Pi(\mathbf{b})]$. Rotation was introduced by Gayler (1998) in order to protect or preserve information because information can be lost/cannot be distinguished because Binding is associative and commutative i.e:- $(\mathbf{a}\otimes\mathbf{b})\otimes(\mathbf{c}\otimes\mathbf{d})=(\mathbf{a}\otimes\mathbf{c})\otimes(\mathbf{b}\otimes\mathbf{d})$. In order to make the make a distinction between $(\mathbf{a}\otimes\mathbf{b})\otimes(\mathbf{c}\otimes\mathbf{d})$ and $(\mathbf{a}\otimes\mathbf{c})\otimes(\mathbf{b}\otimes\mathbf{d})$, rotation is used :- $(\mathbf{a}\otimes\mathbf{b})\otimes\Pi(\mathbf{c}\otimes\mathbf{d})$.
Note that operations and vectors mentioned above have counterparts in the bipolar space $\{-1,1\}^{n}$.
2.4 GNN-based Node Embdeddings
Recall that GNN use the node features, together with the adjacency information to learn the embedding $\mathbf{h}_{v}$ of each node $v\in\mathcal{V}$. Most GNN use a neighborhood aggregation scheme where each node recursively aggregates the features of its neighbors to compute its feature vector Xu et al. (2019); Wu et al. (2020); Zhang et al. (2018); Zhou et al. (2022). After $k$ iterations of such aggregation, the node’s feature vector yields its embedding. More precisely, the latent representation of node $v$ at the $k$th iteration is given by:
$\mathbf{h}^{(k)}_{v}=\text{Agg}(\{\mathbf{W}^{(k)}\mathbf{h}^{(k-1)}_{j};j\in\mathcal{N}^{1}(v)\cup\{v\}\})$
[…]
2.5 Representational Power of GNN
Several authors have investigated the representational power of GNN from different perspectives, including their ability to distinguish between different graph structures Xu et al. (2019), e.g., using the Weisfeiler-Lehman (WL) test Weisfeiler and Leman (1968); Huang and Villar (2021) (which is known to decide graph isomorphism in general, barring a few exceptions, e.g., regular graphs), their ability to learn a set of functions on graphs Scarselli et al. (2009), and the equivalence of the preceding two perspectives Chen et al. (2019). The WL test iteratively updates a node’s feature vector by aggregating the feature vectors of its network neighbors Weisfeiler and Leman (1968); Babai and Kucera (1979). The injectivity of the aggregation function used ensures that different neighborhoods of the nodes map to different feature vectors Huang and Villar (2021). Alternatively, if the set of feature vectors of a node are represented as a multiset, the aggregation function must be able to map different multisets into different representations. As shown by Xu et al. (2019); Morris et al. (2019), GNN are at most as powerful as the WL test in distinguishing graph structures; and for a GNN with a sufficient number of layers will be as powerful as the WL test if the mapping $\mathcal{A}:\mathcal{G}\rightarrow\mathbb{R}^{d}$ produced by the iterative aggregation scheme
$\mathbf{h}^{(k)}_{v}=g(\mathbf{h}^{(k-1)}_{v},f(\{\mathbf{h}^{(k-1)}_{j};j\in\mathcal{N}^{1}(v)\}))$
where $g(\cdot)$ and $f(\cdot)$ and $\mathcal{A}$’s graph-level readout are injective.
GNN generalize the WL test in that, in addition to discriminating different graph structures as they learn to map similar graph structures to similar representations Xu et al. (2019).
3 Hyper-Dimensional Graph Learner
In essence, our design of HDGL aims to achieve a high level of expressive power similar to graph neural networks (GNN). We utilize HD computing operations on node representations, taking advantage of HD operations—Bundling ($\oplus$), Binding ($\otimes$), and Rotation ($\Pi$) —and their properties. This ensures that HDGL, akin to GNN, not only aspires to build expressive power but also learns to map similar graphs to similar representations.
The proposed HD learner carries out 3 major steps:
  1. i
    Map Node features $\mathbf{x}_{v}\in\mathbb{R}^{d}$ to a intermediate HD representation $\mathbf{r}_{v}\in\{0,1\}^{\beta}$ $\forall v\in\mathcal{V}$.
  2. ii
    Construct a latent HD representation $\mathbf{z}_{v}\in\{0,1\}^{\beta}$ using $\mathbf{r}_{v}$ and $\mathcal{N}^{k}(v)$.
  3. iii
    Finally, use $\mathbf{z}_{v}$ to predict $y_{v}$
We proceed to describe each of these steps in detail.
Figure 1: An example mapping node features to a binary space using Random Hyperplanes from Efimov (2023). A point receives a value of 1 when on the left side of a hyperplane and 0 when on the right side. Utilizing multiple hyperplanes results in a binary sketch for each point.
3.1 Mapping Node Features to HD-space
The first crucial step is to map node features from $\mathbb{R}^{d}$ space to $\{0,1\}^{\beta}$ HD space such that $\beta>>d$. We do this using random hyperplane tessellations Dirksen et al. (2022). The key idea is to generate $\beta$ random hyperplanes in $\mathbb{R}^{d}$ and set the corresponding co-ordinate of the HD representation of a node based on whether the its feature vector in $\mathbb{R}^{d}$ lies on the positive or the negative side of the hyperplane. With $\beta$ such hyper-planes in hand, we can generate an $n$ dimensional binary sketch for each node. More precisely, let $\mathbf{A}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{\beta}$ be a matrix with each row $\mathbf{a}_{1}^{T},\cdots\mathbf{a}_{\beta}^{T}$ drawn from $\mathcal{N}(0,I_{d})$ and $\gamma$ be uniformly distributed on $[-\lambda,\lambda]$. For conciseness, we represent $\beta$ samples of $\gamma$ as $\Gamma\sim[-\lambda,\lambda]^{\beta}$. Using $\mathbf{A}$ and $\Gamma$, a randomized binary sketch $\mathbf{r}_{v}$ for a node $v\in\mathcal{V}$ is given by $\mathbf{r}_{v}=\text{sign}(\mathbf{A}\mathbf{x}_{v}+\Gamma)$.
Note that the resulting binary sketch of each node is in fact a $\beta$-dimensional binary HD vector. Nodes with similar features will have similar HD representations; and nodes with dissimilar features will have dissimilar HD representations.
3.2 Neighborhood Aggregation Scheme
Under the simplifying assumption that nodes that have identical multi-sets of features over their $k$-hop neighborhoods are likely to also have matching local topologies, we approximate the injectivity of the embedding $\mathbf{z}_{v}$ which maps each node to its HD representation by aggregating features of nodes over a $k$-hop neighborhood of $v$ as follows:
$\begin{split}\mathbf{z}_{v}=\phi\Big{(}\mathbf{r}_{v},\varphi_{1}(\{\mathbf{r}_{j};j\in\mathcal{N}^{1}(v)\}),\ldots\\ \ldots,\varphi_{K}(\{\mathbf{r}_{m};m\in\mathcal{N}^{K}(v)\})\Big{)}\end{split}$
where $\mathbf{z}_{v}$ denotes the HD embedding of node $v$ and $\phi(\cdot)$ and $\varphi_{k}(\cdot);k\in\{1,\cdots,K\}$ are injective functions.
3.2.1 Bundling for generalizing across similar nodes
We propose that for a node $v\in\mathcal{V}$; the functions $\varphi_{k}(\cdot)$ should be the Bundling operation $\oplus$ since, due to its permutation invariance, allows us to represent unordered sets (or multi-sets). Hence, we propose to use Bundling to realize $\varphi_{k}$. More precisely,
$\varphi_{k}\Big{(}\{\mathbf{r}_{j};j\in\mathcal{N}^{k}(v)\}\Big{)}=\mathop{\oplus}_{j\in\mathcal{N}^{k}(v)}\mathbf{r}_{j}$
where $\mathop{\oplus}_{j\in\mathcal{N}^{k}(v)}\mathbf{r}_{j}$ denotes binding the HD representations of all of the $k$-hop neighbors of $v$. This operation is analogous to finding a high-dimensional vector that is the most similar to all of the input HD-vectors, akin to how we find the mean in Euclidean spaces and like the mean, it is not injective. However, as noted by Xu et al. (2019), GNNs that use the mean to aggregate features over node neighborhoods Kipf and Welling (2017); Hamilton et al. (2017) perform quite well in practice because the mean operation becomes increasingly injective when graphs consist of diverse node features, which is often the case in real-world graphs.
It is worth noting that the Binding operation $\otimes$ would be inappropriate in this context because binding of two similar vectors results in dissimilar vectors. Thus, if Binding were used to realize $\varphi_{k}$, similar yet distinct local neighborhoods structures would yield different HD representations, adversely impacting HDGL’s ability to generalize across such graph structures. To complicate matters, Binding is also invertible i.e $\mathbf{a}\otimes\mathbf{b}\otimes\mathbf{b}=\mathbf{a}$; and hence, it could discard information that is useful for downstream task e.g., node label prediction.
3.2.2 Binding for linking node with its neighborhood
Having incorporated the Bundling operation in our framework by assigning it to the function $\varphi_{k}(\cdot)$, the remaining task involves identifying a suitable operation for the function $\phi(\cdot)$, for coupling the representation of a node obtained using bundling (see above) with the representations of nodes in its neighborhood.
We propose to use Binding the operation $\otimes$ to realize the function $\phi(\cdot)$ because it offers a means to balance the discriminative power of the resulting model against its generalization ability. We illustrate the preceding claim with an example. Let $v\in\mathcal{V}$ and $w\in\mathcal{V}$ be two nodes in a graph. Suppose $\phi(\cdot)$ is realized using the Binding Operation $\otimes$ and $\psi_{k}(\cdot)$ is realized using the Bundling Operation. Considering only 1-hop and 2-hop neighbors of $v$, the latent representation of $v$ and $w$ will be
$\begin{split}\mathbf{z}_{v}&=\mathbf{r}_{v}\otimes\Big{(}\mathop{\oplus}_{j\in\mathcal{N}^{1}(v)}\mathbf{r}_{j}\Big{)}\otimes\Big{(}\mathop{\oplus}_{m\in\mathcal{N}^{2}(v)}\mathbf{r}_{m}\Big{)}\\ \mathbf{z}_{w}&=\mathbf{r}_{w}\otimes\Big{(}\mathop{\oplus}_{j\in\mathcal{N}^{1}(w)}\mathbf{r}_{j}\Big{)}\otimes\Big{(}\mathop{\oplus}_{m\in\mathcal{N}^{2}(w)}\mathbf{r}_{m}\Big{)}\end{split}$ (1)
Suppose we have
$\begin{split}\mathbf{r}_{w}\approx\mathbf{r}_{v};\quad\;\Big{(}\mathop{\oplus}_{j\in\mathcal{N}^{1}(v)}\mathbf{r}_{j}\Big{)}\approx\Big{(}\mathop{\oplus}_{j\in\mathcal{N}^{1}(w)}\mathbf{r}_{j}\Big{)}\\ \end{split}$
and ${\mathop{\oplus}_{j\in\mathcal{N}^{2}(v)}\mathbf{r}_{j}}$ is dissimilar to ${\mathop{\oplus}_{j\in\mathcal{N}^{2}(w)}\mathbf{r}_{j}}$. Then $\mathbf{z}_{v}$ and $\mathbf{z}_{w}$ will be dissimilar since binding is reflecting i.e
$d_{H}\Big{[}\Big{(}\mathop{\oplus}_{j\in\mathcal{N}^{2}(v)}\mathbf{r}_{j}\Big{)},\Big{(}\mathop{\oplus}_{j\in\mathcal{N}^{2}(w)}\mathbf{r}_{j}\Big{)}\Big{]}\approx d_{H}\Big{[}\mathbf{z}_{v},\mathbf{z}_{w}\Big{]}$
It is easy to see that if all the terms in $\mathbf{z}_{v}$ and $\mathbf{z}_{w}$ are similar, i.e., $v$ and $w$ have similar node features, and similar 1-hop and 2-hop neighborhoods, then $\mathbf{z}_{v}\approx\mathbf{z}_{w}$ as desired.
On the other hand, if the terms representing the features of the 2-hop neighbors of both $v$ and $w$ differ, it results in the dissimilarity between $\mathbf{z}_{v}$ and $\mathbf{z}_{w}$ a desired outcome. When not only the 2-hop neighbors but also the 1-hop neighbors of both $v$ and $w$ are dissimilar, the dissimilarity between $\mathbf{z}_{v}$ and $\mathbf{z}_{w}$ becomes even greater. The key property that makes binding attractive for realizing the function $\phi(\cdot)$ is its reflectivity, i.e., the distance between two HD vectors remains unchanged when both of them are bound to the same third vector.
Conversely, it is easy to see that Bundling is not a good choice for realizing $\phi(\cdot)$. Bundling of the node vector $\mathbf{r}_{v}$ with neighborhood information vector $\oplus_{j\in\mathcal{N}^{k}(v)}\mathbf{r}_{j}$, a vector that has already been bundled with $k$-hop neighbors, will result in a vector that predominantly represents information from a node’s neighborhood, while leaving out essential information about the itself. This is due to the fact that Bundling is a bitwise majority operation; and hence, if bundling is used to realize $\phi(\cdot)$, the information from a node’s neighborhood suppresses the information about the node itself.
Dataset Nodes Edges Features Classes Label rate Edge density
CORA 2708 5429 1433 7 0.0563 0.0004
CiteSeer 3327 4732 3703 6 0.0569 0.0004
PubMed 19717 44338 500 3 0.0030 0.0001
Coauthor CS 18333 81894 6805 15 0.0164 0.0001
Coauthor Physics 34493 247962 8415 5 0.0029 0.0001
Table 1: Dataset Summary.
3.2.3 Improving injectivity via Rotation
Because Binding is associative and commutative, it can adversly impact the injectivity and consequently, therepresentational power of our HD model. For example, suppose $v\in\mathcal{V}$ and $w\in\mathcal{V}$ are two nodes in a graph with encoding obtained using HD mapping that uses Bundling to realize $\varphi_{k}$ and binding to realize $\phi(\cdot)$ and $\mathbf{z}_{v}$ and $\mathbf{z}_{w}$ are computed as per Eq. (1). Consider the scenario in which $\mathbf{x}_{v}=\mathbf{x}_{w}$, $\mathcal{N}^{1}(v)=\mathcal{N}^{2}(w)$ and $\mathcal{N}^{2}(v)=\mathcal{N}^{1}(w)$. Now, assuming $\mathcal{N}^{1}(v)\neq\mathcal{N}^{2}(v)$, the representation of $w$ can be rewritten as
$\begin{split}\mathbf{z}_{w}=\mathbf{r}_{v}\otimes\Big{(}\mathop{\oplus}_{j\in\mathcal{N}^{2}(v)}\mathbf{r}_{j}\Big{)}\otimes\Big{(}\mathop{\oplus}_{m\in\mathcal{N}^{1}(v)}\mathbf{r}_{m}\Big{)}\end{split}$
Since Binding is both associative and commutative, the latent representations $\mathbf{z}_{v}$ and $\mathbf{z}_{w}$ become equal, despite their 1-hop and 2-hop neighbors being different. This poses a problem is terms of discriminative power of the model, leading to a lack of expressive power in our HD-algorithm. To address this issue, we utilize the Rotation $\Pi$ operator as it protects/preserves information. To be more specific, with the incorporation of the rotation/permutation operator, the latent representation for nodes $v\in\mathcal{V}$ and $w\in\mathcal{V}$ changes to
$\begin{split}\mathbf{z}_{v}=\mathbf{r}_{v}\otimes\Pi\Big{(}\mathop{\oplus}_{j\in\mathcal{N}^{1}(v)}\mathbf{r}_{j}\Big{)}\otimes\Pi\;\Pi\Big{(}\mathop{\oplus}_{m\in\mathcal{N}^{2}(v)}\mathbf{r}_{m}\Big{)}\\ \mathbf{z}_{w}=\mathbf{r}_{v}\otimes\Pi\Big{(}\mathop{\oplus}_{j\in\mathcal{N}^{2}(v)}\mathbf{r}_{j}\Big{)}\otimes\Pi\;\Pi\Big{(}\mathop{\oplus}_{m\in\mathcal{N}^{1}(v)}\mathbf{r}_{m}\Big{)}\end{split}$
where $\Pi(\cdot)$ and $\Pi\;\Pi(\cdot)$ denote single and double rotation operations respectively. The preceding equation ensures that $\mathbf{z}_{v}$ and $\mathbf{z}_{w}$ will be dissimilar since rotation of a HD vector yields a HD vector that is dissimilar/orthogonal to the original. That is,
$\begin{split}d_{H}\Big{[}\Pi\Big{(}\mathop{\oplus}_{j\in\mathcal{N}^{1}(v)}\mathbf{r}_{j}\Big{)},\Pi\;\Pi\Big{(}\mathop{\oplus}_{m\in\mathcal{N}^{2}(v)}\mathbf{r}_{m}\Big{)}\Big{]}\approx 0.5\\ d_{H}\Big{[}\Pi\Big{(}\mathop{\oplus}_{j\in\mathcal{N}^{2}(v)}\mathbf{r}_{j}\big{)},\Pi\;\Pi\Big{(}\mathop{\oplus}_{m\in\mathcal{N}^{1}(v)}\mathbf{r}_{m}\Big{)}\Big{]}\approx 0.5\\ \end{split}$
Hence, considering only 1-hop and 2-hop neighbors, the proposed aggregation scheme to be used by HDGL for computing the HD representation of a node $v\in\mathcal{V}$ in the graph is given by
$\begin{split}\mathbf{z}_{v}=\mathbf{r}_{v}\otimes\Pi\Big{(}\mathop{\oplus}_{j\in\mathcal{N}^{1}(v)}\mathbf{r}_{j}\Big{)}\otimes\Pi\;\Pi\Big{(}\mathop{\oplus}_{m\in\mathcal{N}^{2}(v)}\mathbf{r}_{m}\Big{)}\\ \end{split}$ (2)
3.3 Node Label Prediction in HDGL
With the procedure by which HDGL constructs the HD representation for each of the nodes in a graph, we turn to the task of using the resulting HD node representation to predict the node labels using $\mathbf{z}_{v}$. Here, we follow the standard approach to predicting class labels of HD encoded objects Ge and Parhi (2020). Specifically, during the learning phase, we Bundle the collection of HD encodings of nodes belonging to each class yielding a HD Class Hyper-Vector for each of the classes. During the inference phase, given the HD vectors for each of the classes, we predict the label of an unlabeled node $u$ as simply the label associated with the class hyper-vector that is, among all class HD class vectors is closest to the HD vector $\mathbf{z}_{u}$, the encoding of $u$. More precisely let $\forall\ell\in\{1,\cdots,L\}$, $\mathcal{T}_{\ell}\subset\mathcal{V}$ be the subset of nodes that make up the training set with label $\ell$ and $\mathcal{V}_{\text{test}}\subset\mathcal{V}$ be the subset of unlabeled nodes in the graph. The class HD are constructed for each class as follows:
$\mathbf{c}_{\ell}=\mathop{\oplus}_{i\in\mathcal{T}_{\ell}}\mathbf{z}_{i}\quad;\;\forall\ell\in\{1,\cdots,L\}$
resulting in a set of class HD vectors $\{\mathbf{c}_{1},\cdots,\mathbf{c}_{L}\}$. Let $u\in\mathcal{V}_{\text{test}}$ be a node for which the class label is to be predicted. We predict the class label for $u$ using its HD representation and the HD representations of each of the classes as follows:
$\hat{y}_{u}=\mathop{\operatorname{argmin}}_{\ell\in\{1,\cdots,L\}}d_{H}[\mathbf{z}_{u},\mathbf{c}_{\ell}].$
4 Experiments
We proceed to report results of extensive experiments comparing HDGl with against a set of state-of-the-art node labeling methods, including several GNN models, in terms of both accuracy and runtime.
4.1 Data sets
For our experiments, we use the three of the most commoly used node labeling benchmarks: CORA McCallum et al. (2000), CiteSeer Giles et al. (1998), PubMed Namata et al. (2012) data sets and two data sets from the the Microsoft Academic Graph: Coauthor CS and Coauthor Physics from Shchur et al. (2018). We used the Train/Validation/Test splits described inYang et al. (2016); Kipf and Welling (2017); Veličković et al. (2018) for CORA, CiteSeer and PubMed data sets and that described in Shchur et al. (2018) for Coauthor CS and Coauthor Physics data sets. The data sets are summarized in Table 1.
4.2 Models
We compare the performance of HDGL with that of several strong baselinesL Multilayer Perceptron (MLP), DeepWalk Perozzi et al. (2014), Label Propagation. We also directly compare HDGL with state-of-the-art models: Graph Convolutional Network (GCN) Kipf and Welling (2017), Graph Attention Network (GAT) Veličković et al. (2018), and lastly, an fast GNN, namely, Simplified Graph Convolutional Network (SGC) Wu et al. (2019) which is essentially a logistic regression model trained using features extracted from $k$-hop neighborhoods of nodes as input.
4.3 HDGL Model Specifications
To ensure that HDGL are directly comparable with most GNN architectures which typically use 2 layers, and hence aggregate information over 2-hop neighbors, we limit HDGL to aggregate information over at most 2-hops i.e HD node representations are calculated as per Eq.(2).
Because Bundling of an even number of HD vectors can result in ties that would need to be broken using tie-breaking policy (See Section 2), we opt to randomly sample an odd number, specifically, 11 of the 1-hop neighbors and 21 of the 2-hop neighbors during the construction of HD node representations by HDGL.
In the case of CORA and CiteSeer datasets, we map the node features to $50,000$-dimensional HD vectors via random projections. For the PubMed, Coauthor CS, and Coauthor Physics datasets, we map the node features $20,000$-dimensional HD vectors in a similar fashion. While these choices may seem arbitrary, we note that a choice between $20000$ and $50000$ is typically large enough for the desired properties of HD-computing to hold.
4.4 Training and Evaluation
Recall that iterative machine learning methods, including GNN typically make use of a validation data set that is distinct from the test data set for deciding when to terminate training, or for tuning the hyper-parameters. Hence, effectively, they make use of both training and validation data sets for building the predictive model and test data for evaluating the trained model’s performance. In contrast, one-shot learning methods such as HDGL have no need for validation data. Hence, to ensure fair comparison between HDGL and other methods, we evaluate their performance on the training data, while letting the individual methods use training and validation data as appropriate for the algorithm in question. Thus, HDGL uses training and validation sets together to build the model, whereas GNN use training data for training the model, validation data for early stopping and hyperparameter tuning. In the case of GNN methods that rely on iterative training, we choose the models with hyperparameters tuned to obtain optimal performance on validation data. For tuning the hyperparameters, we used grid search over a set of 12 configurations. In the case of GCN and GAT, the following hyperparameters were optimized the number of hidden units in the first layer, learning rate, and weight decay. In the case of GCN,networks with 32 and 64 hidden units were explored whereas in the case of GAT, 4 and 8 attention heads, each with 8 hidden units were explored. Both GCN and GAT were explored with weight decays in [1e-3, 1e-4, 5e-4]. In the case of SGC which does not use any hidden layers, only the learning rate and weight decay parameters were tuned in [1e-3, 5e-4, 1e-4, 5e-5, 1e-5, 5e-6]. GCN, GAT, and SGC were tuned for learning rates set to 1e-2 and 1e-3. The early stopping criterion optimized over choices of 25 and 50 epochs, respectively, for all three models. As MLP, DeepWalk, and Label Prop represent well-established baselines, we do not conduct a hyperparameter search. Instead, we choose to utilize publicly released implementations that already incorporate optimal hyperparameters.
All of the models were evaluated on the same test data. In each case, we report measured performance on the test data, averaged over 10 random choices of the relevant parameters: random hyperplanes used to construct HD representations used by HDGL, weight initializations of the deep learning models.
Method Cora Citeseer Pubmed CompSci Physics
MLP 55.1 $\pm$ 0.6 46.5 $\pm$ 0.4 71.4 $\pm$ 0.5 87.9 $\pm$ 0.6 88.2 $\pm$ 1.0
DeepWalk 67.2 $\pm$ 0.8 43.2 $\pm$ 0.6 65.3 $\pm$ 0.5 78.1 $\pm$ 0.9 87.6 $\pm$ 0.9
Label Prop 68.2 $\pm$ 1.1 45.3 $\pm$ 0.8 63.1 $\pm$ 0.5 74.1 $\pm$ 3.9 86.1 $\pm$ 2.1
GCN 81.9 $\pm$ 0.6 70.3 $\pm$ 0.5 79.0 $\pm$ 0.4 91.1 $\pm$ 0.8 92.8 $\pm$ 0.5
GAT 82.8 $\pm$ 0.5 71.0 $\pm$ 0.6 79.1 $\pm$ 0.3 90.5 $\pm$ 0.6 92.5 $\pm$ 0.9
SGC 79.8 $\pm$ 0.0 68.5 $\pm$ 0.1 75.4 $\pm$ 0.0 89.1 $\pm$ 0.1 92.1 $\pm$ 0.1
HDGL 79.5 $\pm$ 0.7 70.7 $\pm$ 0.6 77.8 $\pm$ 1.1 88.9 $\pm$ 0.4 91.0 $\pm$ 0.9
Table 2: Average Test Set Accuracy and Standard Deviation (in Percentage) Over 10 Random Weight Initializations for Deep Learning Models and 10 Randomized Instantiations of HDGL Hyperplanes Across All Datasets
4.5 Implementation Details
All experiments were conducted on Google Colab Pro with High-memory CPU. To ensure fair comparison, although not ideal for HDGL, we used the Python PyTorch framework Paszke et al. (2019), augmented with the Deep Graph Library (DGL) Wang et al. (2019) for importing datasets and implementing deep learning models. None of our experiments used parallel execution or GPUs.
Figure 2: Comparison of Learning Times (in seconds) Across Datasets for HD-GL with GCN, GAT and SGC with Hyperparameter Tuning with Exploration of Search Space with 12 configurations.
4.6 Results
4.6.1 Predictive Performance
The results of our comparison of HDGL with the other methods are shown in Table 2. We observe that HDGL, an one-shot learning method without the need for computationally expensive iterative training or hyperparameter optimization achieves node prediction accuracy that is competitive with those of state-of-the-art GNN models (GCN, GAT, SGC) across all 5 benchmark data sets, in each case reaching performance that is within 1-2% of the best performing model. HDGL also consistently outperforming basic deep learning baseline methods (MLP, DeepWalk, LabelProp).
4.6.2 Run-times
Fig 2 compares run-time of HDGL (which requires no hyperparameter tuning) with those of GCN, GAT, and SGC (with hyper-parameter tuning) with the run-time of HDGL. We observe that barring one exception, PubMed data where the run-times of SGC and HDGL are close, the run-time of HDGL is substantially less than that of all the state-of-the-art GNN.
The speedup obtained by HDGL over the GNN methods is particularly noticeable for larger graph datasets, such as Coauthor CS and Physics. Figure 3, shows how the methods compare with respect to both predictive accuracy achieved and the run-time used to achieve it. It is apparent that HDGL achieves a substantial speedup by factors of 3.5x to 4 with only minimal (1-2%) drop in performance compared to GNN models.
In practice, deep learning models are often optimized using a far more extensive search over the space of hyperparameters, with corresponding increases in run-time, relative to HDGL, with the use of GPUs further widening the gap between GNN and HDGL in terms of their energy footprints.
Figure 3: Tradeoff of Predictive performance against Run-time for HDGL and GNN on Coauthor CS and Coauthor Physics Datasets. Note that GNN models (GCN, GAT, SGC) performance shown is achieved with hyperparameter tuning whereas HDGL offers one-shot learning without the need for hyperparameter tuning
5 Conclusion
We have presented a novel simple, one-shot, and hence fast, and efficient approach to semi-supervised algorithm for graph representation learning. The proposed approach constructs low precision, noise-tolerant, similarity preserving randomized hyper-dimensional representations of nodes in graphs. We take advantage of the close connection between the Weisfeiler-Lehman test for graph isomorphism and a family of graph neural networks to design a hyper-dimensional graph learning algorithm. The proposed algorithm uses hyper-dimensional operators such as bundling and binding to aggregate information from the local neighborhood of each node so as to represent nodes using information that is an approximation of that used by the Weisfeiler-Lehman test.
Results of experiments with widely used benchmark data sets show that the proposed algorithm achieves performance on semi-supervised node labeling tasks that is competitive with the state-of-the-art deep learning methods, without the need for computationally expensive optimization of model parameters and hyper-parameters. Consequently, the proposed method offers an especially attractive alternative with energy footprint that is substantially smaller than that of deep learning based modern graph embedding methods.
Because the model supports one-shot learning of node embeddings, it lends itself to use in settings that call for continual learning algorithms De Lange et al. (2021); Wickramasinghe et al. (2023) wherein data samples, or classes, are presented on an ongoing basis, allowing the model to be continually updated, without the need for retraining on old data. Additional directions for further research include variants of the proposed algorithm that can learn graph representations that invariant or equivariant with respect to a specified set of operations on graphs Satorras et al. (2021); Keriven and Peyré (2019).

### v2

@@
-One-Shot Graph Representation Learning Using Hyperdimensional Computing
+Hyperdimensional Computing for Node Classification and Link Prediction
 Abstract
-We present a novel, simple, fast, and efficient approach for semi-supervised learning on graphs.
-The proposed approach takes advantage of hyper-dimensional computing which encodes data samples using random projections into a high dimensional space (HD space for short).
-Specifically, we propose a Hyper-dimensional Graph Learning (HDGL) algorithm that leverages the injectivity property of the node representations of a family of graph neural networks.
-HDGL maps node features to the HD space and then uses HD operators such as bundling and binding to aggregate information from the local neighborhood of each node.
-Results of experiments with widely used benchmark data sets show that HDGL achieves predictive performance that is competitive with the state-of-the-art deep learning methods, without the need for computationally expensive training.
+We introduce a novel method for transductive learning on graphs using hyperdimensional representations.
+The proposed approach encodes data samples using random projections into a very high-dimensional space (hyperdimensional or HD space for short).
+It obviates the need for expensive iterative training of the sort required by deep learning methods.
+Specifically, we propose a Hyperdimensional Graph Learning (HDGL) algorithm.
+HDGL leverages the injectivity property of node representations of a family of Graph Neural Networks (GNNs) to map node features to the HD space and then uses HD operators such as bundling and binding to aggregate information from the local neighborhood of each node.
+The resulting latent node representations support both node classification and link prediction tasks, unlike typical deep learning methods, which often require separate models for these tasks.
+We report results of experiments using widely used benchmark datasets which demonstrate that, on the node classification task, HDGL is competitive with the state-of-the-art GNN methods with respect to accuracy, at substantially reduced computational cost.
+Furthermore, HDGL is well-suited for class incremental learning where the model has to learn to effectively discriminate between a growing number of classes.
+Our experiments also show that the HD representation constructed by HDGL supports link prediction at accuracies comparable to that of DeepWalk and related methods, although it falls short of state-of-the-art Graph Neural Network (GNN) methods that rely on computationally expensive iterative training.
+We conclude that HDGL offers a computationally efficient alternative to graph neural networks for node classification, especially in settings that call for class-incremental learning or in applications that demand high accuracy models at significantly lower computational cost and learning time than possible with the state-of-the-art graph neural network methods.
 1 Introduction
+Node Classification and Link Prediction in Networks.
 Social networks, citation networks, etc., are naturally represented as graphs where nodes represent the entities that make up the network and edges between pairs of nodes encode relationships between entities.
 For example, in a social network, nodes represent individuals and links denote social connections; in a citation network, nodes represent articles, and (directed) links denote citations.
-Over the past decade, there has been considerable interest in graph learning problems such as node classification and link prediction Li and Pi (2020); Kumar et al.
-(2020).
-Such problems are typically solved using graph embedding techniques Zhou et al.
-(2022); Cai et al.
-(2018); Xu (2021); Wang et al.
-(2022).
-Graph Neural Networks (GNNs), originally introduced by Sperduti and Starita (1997); Frasconi et al.
-(1998); Gori et al.
-(2005); Scarselli et al.
-(2009), which learn low-dimensional embeddings of nodes in a graph by aggregrating information from local neighborhoods of the nodes, have recently emerged as the dominant approach to graph representation learning Kipf and Welling (2017); Veličković et al.
-(2018); Hamilton et al.
-(2017), because of their superior performance across a broad spectrum of applications Wu et al.
-(2020); Zhang et al.
-(2018); Zhou et al.
-(2022).
-Like other deep learning models, GNNs are data hungry, require extensive hyperparameter tuning, and achieve their superior performance at considerable computational cost, resulting in a large carbon footprint with attendant environmental impact Desislavov et al.
-(2023); Strubell et al.
-(2020); de Vries (2023); Thompson et al.
-(2021).
+Over the past decade, there has been considerable interest in graph learning problems such as node classification and link prediction [30, 29].
+Such problems are typically solved using graph embedding techniques [63, 6, 60, 55].
+1.1 Related Work
+Graph Neural Networks (GNNs), originally introduced by Sperduti and Starita [46], Frasconi et al.
+[13], Gori et al.
+[17], Scarselli et al.
+[44], which learn low-dimensional embeddings of nodes in a graph by aggregating information from local neighborhoods of the nodes, have recently emerged as the dominant approach to graph representation learning [26, 53, 18], because of their superior performance across a broad spectrum of applications [58, 62, 63].
+Like other deep learning models, GNNs are data hungry, require extensive hyperparameter tuning, and achieve their superior performance at considerable computational cost, resulting in a large carbon footprint with attendant environmental impact [10, 47, 9, 50].
 Hence, there is an urgent need for computationally more efficient, and consequently, environmentally more sustainable alternatives to deep learning models in general, with a specific emphasis on graph neural networks.
-Hyperdimensional computing Kanerva (1988, 2009); Neubert et al.
-(2019); Thomas et al.
-(2021) offers a brain-inspired alternative to deep neural networks for artificial intelligence and machine learning applications, that maps each object using a sufficiently high-dimensional random encoding (HD encoding for short) to produce a binary or bipolar vector Kanerva (1988, 2009).
-All computations are performed on the objects thus encoded using simple operations, e.g., element-wise additions and dot products Kanerva (1988); Neubert et al.
-(2019); Thomas et al.
-(2021).
-Computations on HD vectors, because they yield binary or bipolar vectors, can be realized using low-precision, fast, low-power, energy-efficient hardware Kleyko et al.
-(2023a, b).
-Hence, there is a growing interest in the use of HD computing in artificial intelligence and machine learning Neubert et al.
-(2019); Kleyko et al.
-(2023b); Ge and Parhi (2020); Morris et al.
-(2022).
-Against this background, we investigate Hyper-dimensional Graph Learner (HDGL), a one-shot approach to graph learning, that makes use of hyper-dimensional computing Kanerva (1988, 2009); Neubert et al.
-(2019); Thomas et al.
-(2021).
-Specifically, we outline a methodology for designing an HD mapping for node embedding that achieves competitive performance in semi-supervised node labeling tasks when compared to GNNs, which necessitate considerably more costly training.
-The key contributions of the paper are as follows:
-  • We show that how to design a HD node embeddings that match the expressive power of GNN embeddings.
-  • We propose HDGL, a graph learner that takes advantage of carefully designed hyperdimensional node embeddings to perform one-shot learning on graphs for semi-supervised node labeling tasks.
+Hyperdimensional computing [22, 23, 37, 49] offers a brain-inspired alternative to deep neural networks for artificial intelligence and machine learning applications, that maps each object using a sufficiently high-dimensional random encoding (HD encoding for short) to produce a binary or bipolar vector [22, 23].
+All computations are performed on the objects thus encoded using simple operations, e.g., element-wise additions and dot products [22, 37, 49].
+Computations on HD vectors, because they yield binary or bipolar vectors, can be realized using low-precision, fast, low-power, energy-efficient hardware [27, 28].
+Hence, there is a growing interest in the use of HD computing in artificial intelligence and machine learning [37, 28, 15, 34].
+Hyperdimensional encodings of graphs have been investigated by Poduval et al.
+[43] and Nunes et al.
+[38].
+These works focus on performing graph operations, e.g., similarity based retrieval, graph classification, etc.
+In contrast, our focus is on methods that learn node-level representations for node classification and link prediction.
+Hyperdimensional node embeddings have been explored by Kang et al.
+[24] and [31].
+Specifically, Kang et al.
+[24] proposed RelHD, a novel processing-in-memory (PIM) hardware architecture [35] based on FeFET technology [4]; and Li et al.
+[31] introduced HyperNode for learning HD embeddings of nodes.
+However, both RelHD and HyperNode suffer from significant shortcomings associated with how they construct node representations from node features (see section 3.5 for details).
+1.2 Overview and Contributions
+We introduce a novel method for transductive learning on graphs using hyperdimensional representations.
+The proposed approach encodes data samples using random projections into a very high- dimensional space (hyperdimensional or HD space for short) [22, 23, 37, 49].
+It obviates the need for expensive iterative training of the sort required by deep learning methods.
+The resulting latent node representations support both node classification and link prediction tasks, unlike typical deep learning methods, which often require separate models for these tasks.
+Contributions. The key contributions of the paper are as follows:
+  • We introduce HDGL, a method for learning HD node embeddings that match the expressive power of GNN embeddings.
+HDGL leverages the injectivity property of node representations of a family of Graph Neural Networks (GNNs) to map node features to the HD space and then uses HD operators such as bundling and binding to aggregate information from the local neighborhood of each node.
+  • We show how the HD node embeddings learned by HDGL can be used to perform both semi-supervised node labeling and link prediction tasks.
+Thus, unlike deep learning methods usually require separate models for each of these tasks, HDGL produces a single model that can support both tasks.
   • We present results of extensive experimental comparisons of HDGL with the state-of-the-art GNN methods on several standard benchmark data sets which show that HDGL achieves performance that is competitive with that of GNN, while avoiding the need for computationally expensive training.
+  • We present results experiments which show that HDGL offers a better alternative to graph neural networks for settings that call for class-incremental learning, where the model is required to effectively discriminate between a growing number of classes that are presented incrementally over time.
+Organization.
 The rest of the paper is organized as follows: Section 3 reviews the HD computing framework, its basic operations, and their properties, and reviews the expressive power of GNN.
-Section 3 introduces HDGL, the proposed HD computing based one-shot algorithm for learning node embeddings.
+Section 4 introduces HDGL, the proposed HD computing based algorithm for learning node embeddings.
 Section 5 describes results of extensive experiments that compare HDGL with the state-of-the-art GNN-based node embedding methods.
 Section 6 concludes with a brief summary of the paper and an outline of some directions for further research.
-2 Background and Preliminaries
-2.1 Node Embedding Problem
+2 Problem Statement
 Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ represents set of undirected nodes, $\mathcal{E}$ represents the set of edges $\mathcal{E}\subseteq\{(u,v)|u,v,\in\mathcal{V}\}$.
@@
 Supposed each node $v\in\mathcal{V}$ is described by a tuple of $d$ features $\mathbf{x}_{v}\in\mathbb{R}^{d}$.
-Let $y_{v}\in\{1,\cdots,L\}$ be the the label associated with each $v\in\mathcal{V}$.
 The $k$-hop neighbors of node $v$ in graph $\mathcal{G}$ are represented using multi-set $\mathcal{N}^{k}(v)$.
-The goal is to learn a latent representation $\phi(v)$ of each node $v$ using the structure of $\mathcal{G}$ and $\mathbf{x}_{v}$ to accurately predict $y_{v}$.
+For the node label prediction problem in the transductive setting, let $y_{w}\in{1,\cdots,L}$ represent the label associated with each node $w\in\mathcal{W}$, where $\mathcal{W}\subset\mathcal{V}$.
+This implies that not all nodes in the graph are labeled.
+The goal is to learn a latent representation $\phi(v)$ for each node $v$, utilizing the structure of $\mathcal{G}$, the feature vector $\mathbf{x}_{v}$, and the labels $y_{w}$, to accurately predict the labels for a set of nodes that are unlabeled in the graph $\mathcal{G}$.In the transductive link prediction setting, the goal is to infer missing edges within the graph by leveraging the latent representations $\phi(v)$ of the nodes.
+3 Background and Preliminaries
 3.1 Properties of HD Computing
-Hyper-dimensional computing, originally introduced by Kanerva (1988) is a brain-inspired approach to cognitive computing that encodes each object $x\in\mathcal{X}$ using a high dimensional static mapping $\phi:\mathcal{X}\rightarrow\mathcal{H}$, where $\mathcal{H}$ is typically $\{-1,1\}^{n}$ or $\{0,1\}^{n}$ where $n$ is sufficiently large Kanerva (1988, 2009).
-All computations are performed in $\mathcal{H}$, using simple operations, e.g., element-wise additions and dot products Kanerva (1988); Neubert et al.
-(2019).
-The mapping $\phi$ is often random, which because its elements are low precision, lends itself to fast, low-power, energy-efficient hardware realizations Kleyko et al.
-(2023a, b).
+Hyper-dimensional computing, originally introduced by Kanerva [22] is a brain-inspired approach to cognitive computing that encodes each object $x\in\mathcal{X}$ using a high dimensional mapping $\phi:\mathcal{X}\rightarrow\mathcal{H}$, where $\mathcal{H}$ is typically $\{-1,1\}^{n}$ or $\{0,1\}^{n}$ where $n$ is sufficiently large [22, 23].
+All computations are performed in $\mathcal{H}$, using simple operations, e.g., element-wise additions and dot products [22, 37].
+The mapping $\phi$ is often random, which because its elements are low precision, lends itself to fast, low-power, energy-efficient hardware realizations [27, 28].
 In this paper, we assume that =$\mathcal{H}=\{0,1\}^{n}$.
@@
 Let $\mathbf{a}\in\{0,1\}^{n}$ and $\mathbf{b}\in\{0,1\}^{n}$ be two such HD vectors where $a_{i}\sim Ber(p=0.5);\forall i\in\{1,\cdots,n\}$ and $b_{i}\sim Ber(p=0.5);\forall i\in\{1,\cdots,n\}$.
-We can then say that $\mathbf{a}$ and $\mathbf{b}$ are near orthogonal or $d_{H}[\mathbf{a},\mathbf{b}]$, the normalized hamming distance between $\mathbf{a}$ and $\mathbf{b}$, $\approx 0.5$ Kanerva (1988).
+We can then say that $\mathbf{a}$ and $\mathbf{b}$ are near orthogonal or $d_{H}[\mathbf{a},\mathbf{b}]$, the normalized hamming distance between $\mathbf{a}$ and $\mathbf{b}$, $\approx 0.5$ [22].
 This is also known as a blessing of HD spaces.
@@
 As noted by Neubert et al.
-(2019), in the case of an $n$-dimensional binary HD space where each of the $n$-dimensional vectors is $f$-sparse, e.g., the number of nonzero entries is $fn$, the resulting HD space has capacity given by $\binom{n}{\lfloor f\cdot n\rfloor}$.
+[37], in the case of an $n$-dimensional binary HD space where each of the $n$-dimensional vectors is $f$-sparse, e.g., the number of nonzero entries is $fn$, the resulting HD space has capacity given by $\binom{n}{\lfloor f\cdot n\rfloor}$.
 For example, when $n=1000$, and $f=0.05$, the resulting HD space can store approximately $10^{80}$ or almost as many patterns as the number of atoms in the observable universe (which is estimated to be between $10^{78}$ and $10^{82}$).
 Noise or Error Tolerance.
-Due to the distributed representation of data across numerous dimensions in HD spaces, errors in HD computing are limited to only a fraction of bits Ahmad and Hawkins (2015); Neubert et al.
-(2019); Thomas et al.
-(2021).
-Thus, HD computing offers noise tolerance without the need for expensive error correction mechanisms Neubert et al.
-(2019).
+Due to the distributed representation of data across numerous dimensions in HD spaces, errors in HD computing are limited to only a fraction of bits [1, 37, 49].
+Thus, HD computing offers noise tolerance without the need for expensive error correction mechanisms [37].
 3.2 HD Computing Operations
@@
 A key property of Rotation is that rotation results in dissimilar/orthogonal vectors i.e $d_{H}[\mathbf{a},\Pi(\mathbf{a})]\approx 0.5$ and it is distributive over both Bundling and Binding and like binding, rotation is also reflective i.e $d_{H}[\mathbf{a},\mathbf{b}]\approx d_{H}[\Pi(\mathbf{a}),\Pi(\mathbf{b})]$.
-Rotation was introduced by Gayler (1998) in order to protect or preserve information because information can be lost/cannot be distinguished because Binding is associative and commutative i.e:- $(\mathbf{a}\otimes\mathbf{b})\otimes(\mathbf{c}\otimes\mathbf{d})=(\mathbf{a}\otimes\mathbf{c})\otimes(\mathbf{b}\otimes\mathbf{d})$.
+Rotation was introduced by Gayler [14] in order to protect or preserve information because information can be lost/cannot be distinguished because Binding is associative and commutative i.e:- $(\mathbf{a}\otimes\mathbf{b})\otimes(\mathbf{c}\otimes\mathbf{d})=(\mathbf{a}\otimes\mathbf{c})\otimes(\mathbf{b}\otimes\mathbf{d})$.
 In order to make the make a distinction between $(\mathbf{a}\otimes\mathbf{b})\otimes(\mathbf{c}\otimes\mathbf{d})$ and $(\mathbf{a}\otimes\mathbf{c})\otimes(\mathbf{b}\otimes\mathbf{d})$, rotation is used :- $(\mathbf{a}\otimes\mathbf{b})\otimes\Pi(\mathbf{c}\otimes\mathbf{d})$.
-Note that operations and vectors mentioned above have counterparts in the bipolar space $\{-1,1\}^{n}$.
+Note that the operations referenced earlier have equivalents in the bipolar space $\{-1,1\}^{n}$.
+Specifically, the Bundle operation can be represented as signed addition in the bipolar space, while the binding operation aligns with multiplication in the bipolar space.
 3.3 GNN-based Node Embdeddings
-Recall that GNN use the node features, together with the adjacency information to learn the embedding $\mathbf{h}_{v}$ of each node $v\in\mathcal{V}$.
-Most GNN use a neighborhood aggregation scheme where each node recursively aggregates the features of its neighbors to compute its feature vector Xu et al.
-(2019); Wu et al.
-(2020); Zhang et al.
-(2018); Zhou et al.
-(2022).
-After $k$ iterations of such aggregation, the node’s feature vector yields its embedding.
-More precisely, the latent representation of node $v$ at the $k$th iteration is given by:
-$\mathbf{h}^{(k)}_{v}=\text{Agg}(\{\mathbf{W}^{(k)}\mathbf{h}^{(k-1)}_{j};j\in\mathcal{N}^{1}(v)\cup\{v\}\})$
 […]
 3.4 Representational Power of GNN
-Several authors have investigated the representational power of GNN from different perspectives, including their ability to distinguish between different graph structures Xu et al.
-(2019), e.g., using the Weisfeiler-Lehman (WL) test Weisfeiler and Leman (1968); Huang and Villar (2021) (which is known to decide graph isomorphism in general, barring a few exceptions, e.g., regular graphs), their ability to learn a set of functions on graphs Scarselli et al.
-(2009), and the equivalence of the preceding two perspectives Chen et al.
-(2019).
-The WL test iteratively updates a node’s feature vector by aggregating the feature vectors of its network neighbors Weisfeiler and Leman (1968); Babai and Kucera (1979).
-The injectivity of the aggregation function used ensures that different neighborhoods of the nodes map to different feature vectors Huang and Villar (2021).
-Alternatively, if the set of feature vectors of a node are represented as a multiset, the aggregation function must be able to map different multisets into different representations.
-As shown by Xu et al.
-(2019); Morris et al.
-(2019), GNN are at most as powerful as the WL test in distinguishing graph structures; and for a GNN with a sufficient number of layers will be as powerful as the WL test if the mapping $\mathcal{A}:\mathcal{G}\rightarrow\mathbb{R}^{d}$ produced by the iterative aggregation scheme
-$\mathbf{h}^{(k)}_{v}=g(\mathbf{h}^{(k-1)}_{v},f(\{\mathbf{h}^{(k-1)}_{j};j\in\mathcal{N}^{1}(v)\}))$
-where $g(\cdot)$ and $f(\cdot)$ and $\mathcal{A}$’s graph-level readout are injective.
-GNN generalize the WL test in that, in addition to discriminating different graph structures as they learn to map similar graph structures to similar representations Xu et al. (2019).
+Several authors have investigated the representational power of GNN, including their ability to distinguish between different graph structures [59], e.g., using the Weisfeiler-Lehman (WL) graph isomorphism test [56, 19] , or to learn a set of functions on graphs [44], or the equivalence of the preceding two notions of expressivity[8].
+The WL test iteratively updates a node’s feature vector by aggregating the feature vectors of its network neighbors [56, 2].
+The injectivity of the aggregation function used ensures that different neighborhoods of the nodes map to different feature vectors [19].
+Alternatively, if the set of feature vectors of a node are represented as a multiset, the aggregation function must be able to map different multisets into distinct encodings.
+As shown by [59, 33], GNN are at most as powerful as the WL test in distinguishing graph structures if the mapping $\mathcal{A}:\mathcal{G}\rightarrow\mathbb{R}^{d}$ produced by the iterative aggregation scheme: $\mathbf{h}^{(k)}_{v}=g(\mathbf{h}^{(k-1)}_{v},f(\{\mathbf{h}^{(k-1)}_{j};j\in\mathcal{N}^{1}(v)\}))$ where $g(\cdot)$ and $f(\cdot)$ and $\mathcal{A}$’s graph-level readout are injective.
+GNN generalize the WL test in that, in addition to discriminating different graph structures as they learn to map similar graph structures to similar representations [59].
+3.5 Limitations of Previous HD Graph Methods
+The HD Computing based graph-level learning methods introduced by Nunes et al.
+[38] and Poduval et al.
+[43] are designed for similarity-based graph retrieval and graph classification, as opposed to node classification or link prediction.
+Nunes et al.
+[38] constructs the Graph HD vector using node features derived using PageRank [5].
+Poduval et al.
+[43] assign random HD-vectors to nodes and considers them as node features to build a Graph HD vector.
+Thus, both methods ignore node features present in the graph.
+RelHD [24] constructs node representations for node classification tasks and uses the resulting representation to design a novel hardware chip design with a scalable processing in-memory (PIM) architecture based FeFET.
+However, RelHD has a significant limitation with respect how it encodes the node features into HD vectors.
+Specifically, it treats node features as dimension-wise indicator variables and computes the HD vector representation of each node by first generating $d$ random basis vector.
+Let $\Upsilon_{p}$ be such a random basis Hyper vector.
+Random basis vectors $\Upsilon_{p}$ are bundled/accumulated if feature at index $p$ is present i.e HD vector for a node $u$ is computed using basis vectors and raw features $\mathbf{x}_{u}=[x_{u1},\ldots,x_{ud}]$ using the following formula:- $\oplus_{p\in\{1,\ldots,d\}}(\mathbbm{1}(x_{up}\neq 0)\cdot\Upsilon_{p})$.
+Consequently, this method is not suitable for graphs with continuous or integer-valued node attributes in continuous form or even simple bag-of-words counts.
+Note that because HyperNode [31] also utilizes the same mechanism for constructing HD vector encodings of the nodes, it suffers from the same limitations as RelHD.
+Furthermore, none of the aforementioned methods [38, 43, 24, 31] support link prediction on graphs.
+In contrast, HDGL offers an effective and efficient method for constructing more expressive node embeddings that can accommodate not only binary or multi-valued, but also integer-valued or continuous node features.
+The resulting HD encodings of nodes support both node classification and link prediction.
 4 Hyper-Dimensional Graph Learner
@@
 This ensures that HDGL, akin to GNN, not only aspires to build expressive power but also learns to map similar graphs to similar representations.
-The proposed HD learner carries out 3 major steps:
-  1. i
+The proposed HD learner carries out 3 major steps: (i).
 Map Node features $\mathbf{x}_{v}\in\mathbb{R}^{d}$ to a intermediate HD representation $\mathbf{r}_{v}\in\{0,1\}^{\beta}$ $\forall v\in\mathcal{V}$.
-  2. ii
+(ii).
 Construct a latent HD representation $\mathbf{z}_{v}\in\{0,1\}^{\beta}$ using $\mathbf{r}_{v}$ and $\mathcal{N}^{k}(v)$.
-  3. iii
-    Finally, use $\mathbf{z}_{v}$ to predict $y_{v}$
+(iii).
+Use $\mathbf{z}_{v}$ to predict $y_{v}$ or predict links/edges in the graph.
 We proceed to describe each of these steps in detail.
-Figure 1: An example mapping node features to a binary space using Random Hyperplanes from Efimov (2023).
+Figure 1: An example mapping node features to a binary space using Random Hyperplanes.
 A point receives a value of 1 when on the left side of a hyperplane and 0 when on the right side.
@@
 The first crucial step is to map node features from $\mathbb{R}^{d}$ space to $\{0,1\}^{\beta}$ HD space such that $\beta>>d$.
-We do this using random hyperplane tessellations Dirksen et al.
-(2022).
+We do this using random hyperplane tessellations [12], a type of Locality Sensitive Hashing [21, 7].
 The key idea is to generate $\beta$ random hyperplanes in $\mathbb{R}^{d}$ and set the corresponding co-ordinate of the HD representation of a node based on whether the its feature vector in $\mathbb{R}^{d}$ lies on the positive or the negative side of the hyperplane.
@@
 Using $\mathbf{A}$ and $\Gamma$, a randomized binary sketch $\mathbf{r}_{v}$ for a node $v\in\mathcal{V}$ is given by $\mathbf{r}_{v}=\text{sign}(\mathbf{A}\mathbf{x}_{v}+\Gamma)$.
+The primary purpose of adding the $\Gamma$ offset is to ensure that the not all sampled hyperplanes pass through the origin.
+When dealing with evenly spread points, typically in the context of real-valued vectors, setting $\Gamma$ to zero is often appropriate.
 Note that the resulting binary sketch of each node is in fact a $\beta$-dimensional binary HD vector.
@@
 4.2 Neighborhood Aggregation Scheme
-Under the simplifying assumption that nodes that have identical multi-sets of features over their $k$-hop neighborhoods are likely to also have matching local topologies, we approximate the injectivity of the embedding $\mathbf{z}_{v}$ which maps each node to its HD representation by aggregating features of nodes over a $k$-hop neighborhood of $v$ as follows:
-$\begin{split}\mathbf{z}_{v}=\phi\Big{(}\mathbf{r}_{v},\varphi_{1}(\{\mathbf{r}_{j};j\in\mathcal{N}^{1}(v)\}),\ldots\\ \ldots,\varphi_{K}(\{\mathbf{r}_{m};m\in\mathcal{N}^{K}(v)\})\Big{)}\end{split}$
-where $\mathbf{z}_{v}$ denotes the HD embedding of node $v$ and $\phi(\cdot)$ and $\varphi_{k}(\cdot);k\in\{1,\cdots,K\}$ are injective functions.
-3.2.1 Bundling for generalizing across similar nodes
+Under the simplifying assumption that nodes that have identical multi-sets of features over their $k$-hop neighborhoods are likely to also have matching local topologies, we approximate the injectivity of the embedding $\mathbf{z}_{v}$ which maps each node to its HD representation by aggregating features of nodes over a $k$-hop neighborhood of $v$ as follows: $\mathbf{z}_{v}=\phi\Big{(}\mathbf{r}_{v},\varphi_{1}(\{\mathbf{r}_{j};j\in\mathcal{N}^{1}(v)\}),\ldots,\varphi_{K}(\{\mathbf{r}_{m};m\in\mathcal{N}^{K}(v)\})\Big{)}$ where $\mathbf{z}_{v}$ denotes the HD embedding of node $v$ and $\phi(\cdot)$ and $\varphi_{k}(\cdot);k\in\{1,\cdots,K\}$ are injective functions.
+Bundling for generalizing across similar nodes.
 We propose that for a node $v\in\mathcal{V}$; the functions $\varphi_{k}(\cdot)$ should be the Bundling operation $\oplus$ since, due to its permutation invariance, allows us to represent unordered sets (or multi-sets).
 Hence, we propose to use Bundling to realize $\varphi_{k}$.
-More precisely,
-$\varphi_{k}\Big{(}\{\mathbf{r}_{j};j\in\mathcal{N}^{k}(v)\}\Big{)}=\mathop{\oplus}_{j\in\mathcal{N}^{k}(v)}\mathbf{r}_{j}$
-where $\mathop{\oplus}_{j\in\mathcal{N}^{k}(v)}\mathbf{r}_{j}$ denotes binding the HD representations of all of the $k$-hop neighbors of $v$.
+More precisely, $\varphi_{k}\Big{(}\{\mathbf{r}_{j};j\in\mathcal{N}^{k}(v)\}\Big{)}=\mathop{\oplus}_{j\in\mathcal{N}^{k}(v)}\mathbf{r}_{j}$ where $\mathop{\oplus}_{j\in\mathcal{N}^{k}(v)}\mathbf{r}_{j}$ denotes binding the HD representations of all of the $k$-hop neighbors of $v$.
 This operation is analogous to finding a high-dimensional vector that is the most similar to all of the input HD-vectors, akin to how we find the mean in Euclidean spaces and like the mean, it is not injective.
-However, as noted by Xu et al.
-(2019), GNNs that use the mean to aggregate features over node neighborhoods Kipf and Welling (2017); Hamilton et al.
-(2017) perform quite well in practice because the mean operation becomes increasingly injective when graphs consist of diverse node features, which is often the case in real-world graphs.
+However, as noted by [59], GNNs that use the mean to aggregate features over node neighborhoods [26, 18] perform quite well in practice because the mean operation becomes increasingly injective when graphs consist of diverse node features, which is often the case in real-world graphs.
 It is worth noting that the Binding operation $\otimes$ would be inappropriate in this context because binding of two similar vectors results in dissimilar vectors.
@@
 To complicate matters, Binding is also invertible i.e $\mathbf{a}\otimes\mathbf{b}\otimes\mathbf{b}=\mathbf{a}$; and hence, it could discard information that is useful for downstream task e.g., node label prediction.
-3.2.2 Binding for linking node with its neighborhood
+Binding for linking node with its neighborhood.
 Having incorporated the Bundling operation in our framework by assigning it to the function $\varphi_{k}(\cdot)$, the remaining task involves identifying a suitable operation for the function $\phi(\cdot)$, for coupling the representation of a node obtained using bundling (see above) with the representations of nodes in its neighborhood.
@@
 It is easy to see that if all the terms in $\mathbf{z}_{v}$ and $\mathbf{z}_{w}$ are similar, i.e., $v$ and $w$ have similar node features, and similar 1-hop and 2-hop neighborhoods, then $\mathbf{z}_{v}\approx\mathbf{z}_{w}$ as desired.
-On the other hand, if the terms representing the features of the 2-hop neighbors of both $v$ and $w$ differ, it results in the dissimilarity between $\mathbf{z}_{v}$ and $\mathbf{z}_{w}$ a desired outcome.
+On the other hand, if the encodings of the features of the 2-hop neighbors of both $v$ and $w$ differ, it results in the dissimilarity between $\mathbf{z}_{v}$ and $\mathbf{z}_{w}$ a desired outcome.
 When not only the 2-hop neighbors but also the 1-hop neighbors of both $v$ and $w$ are dissimilar, the dissimilarity between $\mathbf{z}_{v}$ and $\mathbf{z}_{w}$ becomes even greater.
@@
 This is due to the fact that Bundling is a bitwise majority operation; and hence, if bundling is used to realize $\phi(\cdot)$, the information from a node’s neighborhood suppresses the information about the node itself.
-Dataset Nodes Edges Features Classes Label rate Edge density
-CORA 2708 5429 1433 7 0.0563 0.0004
-CiteSeer 3327 4732 3703 6 0.0569 0.0004
-PubMed 19717 44338 500 3 0.0030 0.0001
-Coauthor CS 18333 81894 6805 15 0.0164 0.0001
-Coauthor Physics 34493 247962 8415 5 0.0029 0.0001
 Table 1: Dataset Summary.
-3.2.3 Improving injectivity via Rotation
+Dataset Nodes Edges Features Feature Type Classes Label Rate Edge Density
+CORA 2708 5429 1433 $\{0,1\}$ 7 0.052 0.0004
+CiteSeer 3327 4732 3703 $\{0,1\}$ 6 0.036 0.0004
+PubMed 19717 44338 500 $(0,1)$ 3 0.003 0.0001
+BlogCatalog 5196 171743 8189 $\mathbb{Z}^{+}$ 6 0.023 0.0127
+Coauthor CS 18333 81894 6805 $\mathbb{Z}^{+}$ 15 0.016 0.0001
+Coauthor Physics 34493 247962 8415 $\mathbb{Z}^{+}$ 5 0.002 0.0001
+DBLP 17725 105781 768 $\mathbb{R}$ 4 0.004 0.0006
+Improving injectivity via Rotation.
 Because Binding is associative and commutative, it can adversly impact the injectivity and consequently, therepresentational power of our HD model.
@@
 Consider the scenario in which $\mathbf{x}_{v}=\mathbf{x}_{w}$, $\mathcal{N}^{1}(v)=\mathcal{N}^{2}(w)$ and $\mathcal{N}^{2}(v)=\mathcal{N}^{1}(w)$.
-Now, assuming $\mathcal{N}^{1}(v)\neq\mathcal{N}^{2}(v)$, the representation of $w$ can be rewritten as
-$\begin{split}\mathbf{z}_{w}=\mathbf{r}_{v}\otimes\Big{(}\mathop{\oplus}_{j\in\mathcal{N}^{2}(v)}\mathbf{r}_{j}\Big{)}\otimes\Big{(}\mathop{\oplus}_{m\in\mathcal{N}^{1}(v)}\mathbf{r}_{m}\Big{)}\end{split}$
+Now, assuming $\mathcal{N}^{1}(v)\neq\mathcal{N}^{2}(v)$, the representation of $w$ can be rewritten as $\mathbf{z}_{w}=\mathbf{r}_{v}\otimes\Big{(}\mathop{\oplus}_{j\in\mathcal{N}^{2}(v)}\mathbf{r}_{j}\Big{)}\otimes\Big{(}\mathop{\oplus}_{m\in\mathcal{N}^{1}(v)}\mathbf{r}_{m}\Big{)}$.
 Since Binding is both associative and commutative, the latent representations $\mathbf{z}_{v}$ and $\mathbf{z}_{w}$ become equal, despite their 1-hop and 2-hop neighbors being different.
@@
 $\begin{split}\mathbf{z}_{v}=\mathbf{r}_{v}\otimes\Pi\Big{(}\mathop{\oplus}_{j\in\mathcal{N}^{1}(v)}\mathbf{r}_{j}\Big{)}\otimes\Pi\;\Pi\Big{(}\mathop{\oplus}_{m\in\mathcal{N}^{2}(v)}\mathbf{r}_{m}\Big{)}\\ \end{split}$ (2)
-3.3 Node Label Prediction in HDGL
+4.3 Node Label Prediction using HDGL
 With the procedure by which HDGL constructs the HD representation for each of the nodes in a graph, we turn to the task of using the resulting HD node representation to predict the node labels using $\mathbf{z}_{v}$.
-Here, we follow the standard approach to predicting class labels of HD encoded objects Ge and Parhi (2020).
+Here, we follow the standard approach to predicting class labels of HD encoded objects [15].
 Specifically, during the learning phase, we Bundle the collection of HD encodings of nodes belonging to each class yielding a HD Class Hyper-Vector for each of the classes.
@@
 More precisely let $\forall\ell\in\{1,\cdots,L\}$, $\mathcal{T}_{\ell}\subset\mathcal{V}$ be the subset of nodes that make up the training set with label $\ell$ and $\mathcal{V}_{\text{test}}\subset\mathcal{V}$ be the subset of unlabeled nodes in the graph.
-The class HD are constructed for each class as follows:
-$\mathbf{c}_{\ell}=\mathop{\oplus}_{i\in\mathcal{T}_{\ell}}\mathbf{z}_{i}\quad;\;\forall\ell\in\{1,\cdots,L\}$
-resulting in a set of class HD vectors $\{\mathbf{c}_{1},\cdots,\mathbf{c}_{L}\}$.
+The class HD are constructed for each class as follows: $\mathbf{c}_{\ell}=\mathop{\oplus}_{i\in\mathcal{T}_{\ell}}\mathbf{z}_{i}\quad;\;\forall\ell\in\{1,\cdots,L\}$, resulting in a set of class HD vectors $\{\mathbf{c}_{1},\cdots,\mathbf{c}_{L}\}$.
 Let $u\in\mathcal{V}_{\text{test}}$ be a node for which the class label is to be predicted.
-We predict the class label for $u$ using its HD representation and the HD representations of each of the classes as follows:
-$\hat{y}_{u}=\mathop{\operatorname{argmin}}_{\ell\in\{1,\cdots,L\}}d_{H}[\mathbf{z}_{u},\mathbf{c}_{\ell}].$
+We predict the class label for $u$ using its HD representation and the HD representations of each of the classes as follows: $\hat{y}_{u}=\mathop{\operatorname{argmin}}_{\ell\in\{1,\cdots,L\}}d_{H}[\mathbf{z}_{u},\mathbf{c}_{\ell}]$.
+4.4 Link Prediction Using HDGL
+Section 4.1 and 4.2 outline a technique for constructing node representations optimized for node classification.
+Here we show that the same node representation can be used to support the transductive link prediction task.
+Kanerva [22, 23] describes a method for encoding entries of a dataset.
+Let $\mathbf{a}_{1}$, $\mathbf{a}_{2}$, $\mathbf{a}_{3}$ be some randomly sampled HD vectors representing data entry attributes (eg.
+weight, height and gender).
+Let $\mathbf{\alpha}$, $\mathbf{\beta}$ and $\mathbf{\gamma}$ be values for the corresponding attributes represented in HD format.
+Let $\mathbf{r}$ be a HD vector such that: $\mathbf{r}=(\mathbf{a}_{1}\otimes\mathbf{\alpha})\oplus(\mathbf{a}_{2}\otimes\mathbf{\beta})\oplus(\mathbf{a}_{3}\otimes\mathbf{\gamma}).$ From the properties of HD-computing [22], it is possible to retrieve values for a corresponding attribute using the Binding $\otimes$ operator.
+Specifically, $\mathbf{\alpha}\approx\mathbf{r}\otimes\mathbf{a}_{1};\quad\mathbf{\beta}\approx\mathbf{r}\otimes\mathbf{a}_{2};\quad\mathbf{\gamma}\approx\mathbf{r}\otimes\mathbf{a}_{3};$
+We apply the same concept to link prediction.
+Let $\mathbf{e}^{+}$ be a hypervector storing information about all or some of the edges present in the graph, and let $\mathbf{e}^{-}$ be a hypervector representing information about some of the non-edges in the graph.
+Formally: $\mathbf{e}^{+}=\mathop{\oplus}_{(u,v)\in\mathcal{E}}(\mathbf{z}_{u}\otimes\mathbf{z}_{v});\quad\mathbf{e}^{-}=\mathop{\oplus}_{(u,v)\notin\mathcal{E}}(\mathbf{z}_{u}\otimes\mathbf{z}_{v}).$ Let $\mathbf{Z}$ have $N$ rows, storing latent representations of all the nodes in the graph.
+That is, a row in $\mathbf{Z}$ is represented by $\mathbf{z}_{u}$, where $u\in\mathcal{V}$.
+Now that we have encoded edges and non-edges in the graph, the goal is to find new or missing links in the graph.
+This can also be thought of as reconstructing/generating the adjacency matrix.
+To find possible new links we Bind $\otimes$ the edge information hypervector $\mathbf{e}^{+}$ and the latent representation of all the nodes in the graph.
+This results in a new matrix $\mathbf{Z}^{+}$ where each row is $\mathbf{z}_{u}\otimes\mathbf{e}^{+}$ where $u\in\mathcal{V}$.
+$\mathbf{Z}^{+}$ represents the possible node representations of the target nodes in a link, with the corresponding source node representation being $\mathbf{Z}$.
+We follow the same procedure for the non-edge hypervector, resulting in $\mathbf{Z}^{-}$.
+Precisely:
+$\mathbf{Z}^{+}$ $=\begin{bmatrix}\mathbf{z}_{1}\otimes\mathbf{e}^{+}\\ \vdots\\ \mathbf{z}_{N}\otimes\mathbf{e}^{+}\end{bmatrix}$ $\mathbf{Z}^{-}$ $=\begin{bmatrix}\mathbf{z}_{1}\otimes\mathbf{e}^{-}\\ \vdots\\ \mathbf{z}_{N}\otimes\mathbf{e}^{-}\end{bmatrix}$
+Now, we find pairwise normalized distances between rows of $\mathbf{Z}^{+}$ and rows of $\mathbf{Z}$.
+This results in $\mathbf{D}^{+}=dist(\mathbf{Z}^{+},\mathbf{Z})\in[0,1]^{N\times N}$ , a matrix representing distances.
+If entry $(i,j)$ is close to zero, it indicates that the distance between row $i$ of $\mathbf{Z}^{+}$ is close to row $j$ in $\mathbf{Z}$, suggesting that $(i,j)$ might be a possible link.
+We perform the same process for $\mathbf{Z}^{-}$, resulting in the $\mathbf{D}^{-}=dist(\mathbf{Z}^{-},\mathbf{Z})\in[0,1]^{N\times N}$ matrix.
+However, an entry close to zero in $\mathbf{D}^{-}$ indicates a high possibility of not having an edge between the corresponding nodes.
+We now construct a prediction of the adjacency matrix represented by $\hat{\mathbf{A}}$, where entries $\hat{A}_{ij}\in[0,1]$.
+To construct the final reconstructed or predicted adjacency matrix of probabilities (based on distances), which factors in both the distances of having, or not having an edge using $\mathbf{Z}^{+}$ and $\mathbf{Z}^{-}$, we compute:
+$\hat{A}_{ij}=\begin{cases}\sigma\big{(}(1-D^{+}_{ij})+D^{-}_{ij}\big{)}&\text{if }D^{+}_{ij}<D^{-}_{ij}\\ \sigma\big{(}D^{+}_{ij}-(1-D^{-}_{ij})\big{)}&\text{else}\end{cases}$
+where $\sigma(\cdot)$ is the sigmoid function.
+The rationale behind this operation lies in the comparison between $D^{+}_{ij}$ and $D^{-}_{ij}$.
+When $D^{+}_{ij}<D^{-}_{ij}$, it indicates a higher likelihood of a connection between nodes $i$ and $j$ according to the distance metric.
+Hence, we apply the operation $\sigma((1-D^{+}_{ij})+D^{-}_{ij})$ to obtain a value closer to one.
+We use the same logic when $D^{+}_{ij}\geq D^{-}_{ij}$ to obtain a value close to zero.
 5 Experiments
-We proceed to report results of extensive experiments comparing HDGl with against a set of state-of-the-art node labeling methods, including several GNN models, in terms of both accuracy and runtime.
+We proceed to report results of extensive experiments comparing HDGL with a set of state-of-the-art node classification methods and link prediction methods, including several GNN models, in terms of both predictive performance and runtime.
 5.1 Data sets
-For our experiments, we use the three of the most commoly used node labeling benchmarks: CORA McCallum et al.
-(2000), CiteSeer Giles et al.
-(1998), PubMed Namata et al.
-(2012) data sets and two data sets from the the Microsoft Academic Graph: Coauthor CS and Coauthor Physics from Shchur et al.
-(2018).
-We used the Train/Validation/Test splits described inYang et al.
-(2016); Kipf and Welling (2017); Veličković et al.
-(2018) for CORA, CiteSeer and PubMed data sets and that described in Shchur et al.
-(2018) for Coauthor CS and Coauthor Physics data sets.
-The data sets are summarized in Table 1.
+Since HDGL is designed in such a way that nodes with similar features and neighborhoods will have similar representations, we exclusively concentrate on homophilous graphs in this study, deferring the exploration of more complex cases involving heterophily graphs to future research.
+For our experiments, we utilize commonly used node labeling homophilous graph benchmarks: the CORA graph dataset [61], sourced from McCallum et al.
+[32]; the CiteSeer graph dataset [61], sourced from Giles et al.
+[16]; and the PubMed dataset [36].
+Additionally, we include three more datasets from the Microsoft Academic Graph: Coauthor CS and Coauthor Physics from Shchur et al.
+[45], and the BlogCatalog Graph dataset from Huang et al.
+[20].
+Additionally, we introduce a new version of the DBLP dataset, built upon the works of Pan et al.
+[39].
+We remove nodes with no neighbors in the graph obtained from Pan et al.
+[39].
+The original node features from Pan et al.
+[39] consist of manuscript titles.
+We utilize BERT [11] to obtain embeddings for the titles and employ them as node features.
+The data sets and their statistics are given in Table 1.
+We adhere to the Train/Validation/Test splits described in Yang et al.
+[61], Kipf and Welling [26], and Veličković et al.
+[53] for the CORA, CiteSeer, and PubMed datasets, respectively.
+For the Coauthor CS, Coauthor Physics, BlogCatalog, and DBLP datasets, we employ the data splitting strategy outlined in Shchur et al.
+[45].
+We test link prediction models on CORA, CiteSeer, and Pubmed datasets.
+These models are trained on incomplete data, retaining node features but omitting some edges.
+Validation and test sets include excluded edges and randomly selected unconnected node pairs.
+Evaluation assesses the models’ ability to classify both edges and non-edges.
+We follow splitting strategy from Kipf and Welling [25] work i.e Validation and test sets each contain 5% and 10% of citation links.
 5.2 Models
-We compare the performance of HDGL with that of several strong baselinesL Multilayer Perceptron (MLP), DeepWalk Perozzi et al.
-(2014), Label Propagation.
-We also directly compare HDGL with state-of-the-art models: Graph Convolutional Network (GCN) Kipf and Welling (2017), Graph Attention Network (GAT) Veličković et al.
-(2018), and lastly, an fast GNN, namely, Simplified Graph Convolutional Network (SGC) Wu et al.
-(2019) which is essentially a logistic regression model trained using features extracted from $k$-hop neighborhoods of nodes as input.
+For semi-supervised node classifcation task, we compare the performance of HDGL with that of several strong baselines Logistic Regression (LogReg), DeepWalk [42] and Label Propagation [64].
+We also directly compare HDGL with state-of-the-art models: Graph Convolutional Network (GCN) [26], Graph Attention Network (GAT) [53], and lastly, a fast GNN model, namely, Simplified Graph Convolutional Network (SGC) [57] which is essentially a logistic regression model trained using features extracted from $k$-hop neighborhoods of nodes as input.
+We also assess our model against RelHD [24], an existing HD algorithm used for node classification.
+For the transductive link prediction task, we evaluate our model against standard link prediction baselines:- DeepWalk [42] and Spectral clustering (SC) [48].
+We also compare our model against graph autoencoder models:- Graph Autoencoder (GAE) and Variational Graph Autoencoder (VGAE) from Kipf and Welling [25].
 5.3 HDGL Model Specifications
 To ensure that HDGL are directly comparable with most GNN architectures which typically use 2 layers, and hence aggregate information over 2-hop neighbors, we limit HDGL to aggregate information over at most 2-hops i.e HD node representations are calculated as per Eq.(2).
-Because Bundling of an even number of HD vectors can result in ties that would need to be broken using tie-breaking policy (See Section 2), we opt to randomly sample an odd number, specifically, 11 of the 1-hop neighbors and 21 of the 2-hop neighbors during the construction of HD node representations by HDGL.
-In the case of CORA and CiteSeer datasets, we map the node features to $50,000$-dimensional HD vectors via random projections.
-For the PubMed, Coauthor CS, and Coauthor Physics datasets, we map the node features $20,000$-dimensional HD vectors in a similar fashion.
-While these choices may seem arbitrary, we note that a choice between $20000$ and $50000$ is typically large enough for the desired properties of HD-computing to hold.
+Because Bundling of an even number of HD vectors can result in ties that would need to be broken using tie-breaking policy (See Section 3.2), we opt to randomly sample an odd number, specifically, 11 of the 1-hop neighbors and 21 of the 2-hop neighbors during the construction of HD node representations by HDGL for datasets except BlogCatalog.
+For BlogCatalog we sample 21 of the 1-hop neighbors and 51 of the 2-hop neighbors since Blogcatalog graph is denser than the other datasets.
+For node classification in CORA, CiteSeer, and BlogCatalog datasets, we map node features to $50,000$-dimensional HD vectors through random projections.
+Similarly, for the PubMed, Coauthor CS, Coauthor Physics, and DBLP datasets, we map node features to $20,000$-dimensional HD vectors using the same method.
+While these choices may seem arbitrary, it’s crucial to note that selecting dimensions between $20,000$ and $50,000$ typically suffices to maintain the desired properties of HD-computing.
+This assertion finds extensive support in the research from Kanerva [22, 23].
+Table 2: Average Test Set Accuracy and Standard Deviation (in Percentage) Over 10 Random Weight Initializations for Deep Learning Models and 10 Randomized Instantiations of RelHD and HDGL Across Datasets (* denotes datasets with $\{0,1\}$ features.
+In these datasets, RelHD performs comparably to HDGL)
+Method Cora Citeseer Pubmed BlogCat CompSci Physics DBLP
+LogReg 52.2 $\pm$ 0.5 46.3 $\pm$ 0.4 67.9 $\pm$ 0.5 65.3 $\pm$ 1.0 86.8 $\pm$ 0.6 88.2 $\pm$ 0.9 56.2 $\pm$ 3.0
+DeepWalk 65.3 $\pm$ 1.2 46.2 $\pm$ 0.9 69.3 $\pm$ 0.9 58.6 $\pm$ 1.3 78.1 $\pm$ 0.9 86.4 $\pm$ 0.9 66.1 $\pm$ 2.2
+Label Prop 70.9 $\pm$ 0.0 47.7 $\pm$ 0.0 71.3 $\pm$ 0.0 47.8 $\pm$ 3.8 76.6 $\pm$ 0.9 85.6 $\pm$ 1.1 66.4 $\pm$ 1.6
+GCN 81.9 $\pm$ 0.6 70.3 $\pm$ 0.5 79.0 $\pm$ 0.4 69.5 $\pm$ 0.4 91.0 $\pm$ 0.7 92.3 $\pm$ 0.6 72.1 $\pm$ 0.8
+GAT 82.8 $\pm$ 0.5 72.0 $\pm$ 0.6 79.1 $\pm$ 0.3 65.3 $\pm$ 1.2 90.9 $\pm$ 0.6 92.5 $\pm$ 0.9 75.4 $\pm$ 1.0
+SGC 79.8 $\pm$ 0.5 68.5 $\pm$ 0.4 75.4 $\pm$ 0.5 70.8 $\pm$ 0.3 89.1 $\pm$ 0.6 92.1 $\pm$ 0.7 76.3 $\pm$ 0.9
+RelHD 78.9 $\pm$ 0.5 69.3 $\pm$ 0.4 69.1 $\pm$ 0.4 60.9 $\pm$ 1.2 85.5 $\pm$ 0.8 85.3 $\pm$ 0.9 44.9 $\pm$ 0.2
+HDGL 79.5 $\pm$ 0.7 70.0 $\pm$ 0.6 76.8 $\pm$ 1.1 69.3 $\pm$ 0.5 88.9 $\pm$ 0.4 91.0 $\pm$ 0.9 68.7 $\pm$ 1.2
 5.4 Training and Evaluation
@@
 Hence, effectively, they make use of both training and validation data sets for building the predictive model and test data for evaluating the trained model’s performance.
-In contrast, one-shot learning methods such as HDGL have no need for validation data.
+In contrast, learning methods such as HDGL which invole no iterative training do not need validation data.
 Hence, to ensure fair comparison between HDGL and other methods, we evaluate their performance on the training data, while letting the individual methods use training and validation data as appropriate for the algorithm in question.
-Thus, HDGL uses training and validation sets together to build the model, whereas GNN use training data for training the model, validation data for early stopping and hyperparameter tuning.
+Thus, HDGL uses training and validation sets together for learning, whereas GNN use training data for training the model, validation data for early stopping and hyperparameter tuning.
 In the case of GNN methods that rely on iterative training, we choose the models with hyperparameters tuned to obtain optimal performance on validation data.
@@
 GCN, GAT, and SGC were tuned for learning rates set to 1e-2 and 1e-3.
-The early stopping criterion optimized over choices of 25 and 50 epochs, respectively, for all three models.
-As MLP, DeepWalk, and Label Prop represent well-established baselines, we do not conduct a hyperparameter search.
-Instead, we choose to utilize publicly released implementations that already incorporate optimal hyperparameters.
-All of the models were evaluated on the same test data.
-In each case, we report measured performance on the test data, averaged over 10 random choices of the relevant parameters: random hyperplanes used to construct HD representations used by HDGL, weight initializations of the deep learning models.
-Method Cora Citeseer Pubmed CompSci Physics
-MLP 55.1 $\pm$ 0.6 46.5 $\pm$ 0.4 71.4 $\pm$ 0.5 87.9 $\pm$ 0.6 88.2 $\pm$ 1.0
-DeepWalk 67.2 $\pm$ 0.8 43.2 $\pm$ 0.6 65.3 $\pm$ 0.5 78.1 $\pm$ 0.9 87.6 $\pm$ 0.9
-Label Prop 68.2 $\pm$ 1.1 45.3 $\pm$ 0.8 63.1 $\pm$ 0.5 74.1 $\pm$ 3.9 86.1 $\pm$ 2.1
-GCN 81.9 $\pm$ 0.6 70.3 $\pm$ 0.5 79.0 $\pm$ 0.4 91.1 $\pm$ 0.8 92.8 $\pm$ 0.5
-GAT 82.8 $\pm$ 0.5 71.0 $\pm$ 0.6 79.1 $\pm$ 0.3 90.5 $\pm$ 0.6 92.5 $\pm$ 0.9
-SGC 79.8 $\pm$ 0.0 68.5 $\pm$ 0.1 75.4 $\pm$ 0.0 89.1 $\pm$ 0.1 92.1 $\pm$ 0.1
-HDGL 79.5 $\pm$ 0.7 70.7 $\pm$ 0.6 77.8 $\pm$ 1.1 88.9 $\pm$ 0.4 91.0 $\pm$ 0.9
-Table 2: Average Test Set Accuracy and Standard Deviation (in Percentage) Over 10 Random Weight Initializations for Deep Learning Models and 10 Randomized Instantiations of HDGL Hyperplanes Across All Datasets
+The early stopping criterion optimized over choices of 25 and 50 epochs, respectively, for all three GNN models.
+As LogReg, DeepWalk, and Label Prop represent well-established baselines, we do not conduct a hyperparameter search.
+For Label Propagation, we use 20 iterations of propagation (except for blogcatalog dataset where we use 10) and set the alpha factor  [^1] Alpha factor in Label Propagation determines the balance between existing and propagated labels.
+Higher alpha prioritizes existing labels; lower alpha favors propagated ones.
+to 0.5.
+For DeepWalk, we utilize the DGL library, setting the embedding dimension to 32 and the number of walks to 20, with a batch size of 64.
+We sample such walks for all nodes in the graph and train DeepWalk model for 10 epochs and then employ logistic regression for classification, incorporating early stopping based on the validation set.
+Following Kang et al. [24], the RelHD algorithm was employed with 10,000 dimensional random basis HD vectors, as the performance of RelHD plateaus around this dimensionality.
+We report measured performance on the test data, averaged across 10 random selections of pertinent parameters.
+These parameters include the random hyperplanes employed for constructing high-dimensional representations used by HDGL, random basis vector initializations for RelHD, weight initializations of the deep learning models, and data splits for datasets such as Coauthor Physics, Coauthor CS, BlogCatalog, and DBLP.
+For Cora, Citeseer, and Pubmed, we use predefined splits given in Yang et al.
+[61].
+For the link prediction task, we rely on the results directly obtained from the paper for GVAE from Kipf and Welling [25].
 5.5 Implementation Details
 All experiments were conducted on Google Colab Pro with High-memory CPU.
-To ensure fair comparison, although not ideal for HDGL, we used the Python PyTorch framework Paszke et al.
-(2019), augmented with the Deep Graph Library (DGL) Wang et al.
-(2019) for importing datasets and implementing deep learning models.
+To ensure fair comparison, although not ideal for HDGL, we used the Python PyTorch framework [40], complemented by the Deep Graph Library (DGL) [54] for importing datasets and implementing deep learning models.
 None of our experiments used parallel execution or GPUs.
-Figure 2: Comparison of Learning Times (in seconds) Across Datasets for HD-GL with GCN, GAT and SGC with Hyperparameter Tuning with Exploration of Search Space with 12 configurations.
+We have also developed a PyTorch version of RelHD, as the codebase was not publicly available and experiments detailed in Kang et al.
+[24] were reliant on hardware architecture implementations.
+We undertake this effort to compare with our model and use RelHD as a baseline in this study.
 5.6 Results
-4.6.1 Predictive Performance
+Node Classification Performance.
 The results of our comparison of HDGL with the other methods are shown in Table 2.
-We observe that HDGL, an one-shot learning method without the need for computationally expensive iterative training or hyperparameter optimization achieves node prediction accuracy that is competitive with those of state-of-the-art GNN models (GCN, GAT, SGC) across all 5 benchmark data sets, in each case reaching performance that is within 1-2% of the best performing model.
-HDGL also consistently outperforming basic deep learning baseline methods (MLP, DeepWalk, LabelProp).
-4.6.2 Run-times
-Fig 2 compares run-time of HDGL (which requires no hyperparameter tuning) with those of GCN, GAT, and SGC (with hyper-parameter tuning) with the run-time of HDGL.
-We observe that barring one exception, PubMed data where the run-times of SGC and HDGL are close, the run-time of HDGL is substantially less than that of all the state-of-the-art GNN.
-The speedup obtained by HDGL over the GNN methods is particularly noticeable for larger graph datasets, such as Coauthor CS and Physics.
-Figure 3, shows how the methods compare with respect to both predictive accuracy achieved and the run-time used to achieve it.
-It is apparent that HDGL achieves a substantial speedup by factors of 3.5x to 4 with only minimal (1-2%) drop in performance compared to GNN models.
-In practice, deep learning models are often optimized using a far more extensive search over the space of hyperparameters, with corresponding increases in run-time, relative to HDGL, with the use of GPUs further widening the gap between GNN and HDGL in terms of their energy footprints.
-Figure 3: Tradeoff of Predictive performance against Run-time for HDGL and GNN on Coauthor CS and Coauthor Physics Datasets.
-Note that GNN models (GCN, GAT, SGC) performance shown is achieved with hyperparameter tuning whereas HDGL offers one-shot learning without the need for hyperparameter tuning
+We observe that HDGL without the need for computationally expensive iterative training or hyperparameter optimization achieves node prediction accuracy that is competitive with those of state-of-the-art GNN models (GCN, GAT, SGC) across most of the benchmark data sets, in each case reaching performance that is within 1-2% of the best performing model.
+HDGL also consistently outperforming basic deep learning baseline methods (LogReg, DeepWalk, LabelProp).
+The notable advantage of HDGL over RelHD lies in its ability to accommodate features beyond the binary $\{0,1\}$ range, as evident from the results presented in Table 2, where HDGL significantly outperforms.
+This is particularly notable with datasets lacking $\{0,1\}$ features.
+However, when dealing with datasets featuring $\{0,1\}$ attributes like CORA and Citeseer, the performance of HDGL and RelHD appears nearly identical.
+To highlight the advantage of HDGL more prominently, we conducted experiments for node classification on the CORA and CiteSeer datasets where we reduced the node features to 100 dimensions using PCA [41], thereby trying to enforce continuous node features.
+As depicted in Figure 2, the impact is evident through a noticeable drop in performance for RelHD.
+Conversely, for other methods including HDGL, the performance remains relatively unchanged.
+Figure 2: Impact of performing PCA with a dimensionality of 100 on features for node classification accuracy.
+Figure 3: Comparison of Learning Times (in seconds) i.e Across Datasets for HDGL and RelHD with GCN, GAT and SGC with Hyperparameter Tuning with Exploration of Search Space with 12 configurations.
+Runtime for HDGL and RelHD exclusively includes only time Elapsed for learning label representations using train/validation data.
+Table 3: Link prediction results averaged over 10 Randomized Instantiations of HDGL and datasplits. Results for other models are from Kipf and Welling [25] paper.
+Methods Cora Citeseer PubMed
+AUC AP AUC AP AUC AP
+DeepWalk 0.831 0.850 0.805 0.836 0.844 0.841
+SC 0.846 0.885 0.791 0.826 0.849 0.888
+GAE 0.910 0.920 0.895 0.899 0.964 0.965
+VGAE 0.914 0.926 0.908 0.920 0.944 0.947
+HDGL 0.849 0.880 0.768 0.842 0.853 0.884
+Node Classification Run-times.
+Figure 3 compares the run-time for the node classification task between HDGL and RelHD (both of which requires no hyperparameter tuning) with those of GCN, GAT, and SGC (with hyperparameter tuning), alongside the run-time of HDGL.
+We observe that, the run-time of HDGL is substantially less than that of all the state-of-the-art GNNs.
+While in most cases, RelHD exhibits slightly or significantly worse run times compared to HDGL, it’s crucial to note that HDGL consistently delivers superior performance across various datasets when compared to RelHD.
+Table 4: Class Incremental Node Labeling Results.
+No hyperparameter grid search is used for GCN and SGC algorithms.
+Wall-time includes both training and inference of the model.
+HDGL does not require retraining from $t=2$.
+Model BlogCatalog Coauthor Physics
+t = 1 t = 2 t = 3 t = 4 t = 5 t=1 t=2 t=3 t=4
+GCN Accuracy 94.94 85.81 75.16 74.82 69.61 98.95 97.33 95.57 92.58
+Runtime 51.2s 18.7s 16.5s 33.9s 23.9s 65.9s 62.3s 65.8s 51.2s
+SGC Accuracy 94.56 82.31 73.23 72.72 69.81 98.94 97.39 95.36 91.81
+Runtime 6.3s 3.72s 3.43s 3.73s 3.6s 9.1s 8.5s 8.8s 9.7s
+HDGL Accuracy 89.11 88.49 75.56 71.73 69.91 98.56 94.87 92.53 90.8
+Runtime 20.2s 0.3s 0.4s 0.4s 0.6s 38.3s 1.5s 1.7s 1.7s
+It is crucial to note that while we have used grid search on 12 potential setups for the deep learning model hyperparameter optimization of deep learning models typically explores a much larger space.
+This expanded search naturally comes with increased runtime, especially when GPUs are utilized, further widening the energy efficiency gap between Graph Neural Networks (GNN) and traditional deep learning models like HDGL.
+Link Prediction Performance.
+Even though HDGL representations are primarily tailored for transductive node classification, their performance for link prediction is comparable to that of DeepWalk and Spectral clustering; as seen from Table 3.
+However, they do not reach the same level of effectiveness as GAE and VGAE, which leverage GNN methods.
+Despite lacking iterative training like deep learning methods, HDGL still manages to achieve respectable results in link prediction.
+Moreover, unlike deep learning approaches that typically require distinct models for link and node classification, HDGL tackles both tasks within a unified framework.
+We also perform link prediction experiments by varying the dimensionality of the HDGL model.
+This exploration arises from observing a slight yet discernible variation in performance for link prediction tasks, unlike the more consistent outcomes observed in node classification experiments.
+As mentioned earlier, typically, dimensions ranging from 20,000 to 50,000 are sufficient for HD properties to hold.
+However, higher dimensionality generally leads to better performance.
+Nevertheless, beyond a certain threshold, the performance tends to plateau, as illustrated in Figure 4.
+Figure 4: Link Prediction Performance of HDGL under Various dimensionality configurations.
+HDGL in the Class-Incremental Learning Setting Consider a scenario with a dynamic graph where nodes and edges remain static, but the new node labels appear over time i.e initially, only a subset of nodes is labeled, belonging to two classes.
+Over time, new nodes receive labels; thus, introducing new classes.
+This is a node labeling instance of the class-incremental learning problem [52, 3, 51].
+In this setting, deep learning methods typically require either retraining on the data for previously learned classes along with the data for the new classes to avoid catastrophic forgetting of the previously learned classes [51].
+This retraining incurs significant additional runtime and computation.
+In contrast to deep neural networks, which necessitate retraining, Hyperdimensional Graph Learner (HDGL) operates without an iterative training phase.
+Instead, it constructs label hypervectors using predefined latent node representations.
+This approach eliminates the need for retraining, resulting in substantial savings in both runtime and computation costs.
+To simulate incremental node classification on a fully labeled graph, we use BlogCatalog and Coauthor Physics Dataset.
+We first split the nodes into train/val/test as in Shchur et al.
+[45].
+We mask the train/val/test such that at time $t=1$ only the labels $\{1,2\}$ are available; and at $t=2$ only the labels $\{1,2,3\}$ are only available; and so on such that at the final time step, all of the labels $\{1,\cdots,L\}$ are available.
+To simplify matters for GNN models, we assume that the models are provided the list of $L$ possible labels although only a subset have been encountered in the training data.
+This allows the last layer of the GNN to have $L$ outputs.
+Also, we forego hyperpameter search for the GNN models to minimize the computational overhead.
+In comparing HDGL against GCN and SGC in our experimental setup, we opted not to include the GAT model due to its longer runtime, as depicted in Figure 3.
+As seen from Table 4, we see that HDGL incurs a significant computational overhead only at timestep 1, necessitating the computation of node representations followed by inference.
+However, for subsequent timesteps , HDGL leverages previously computed node representations, requiring only inference.
+In contrast, deep learning models like GCN and SGC mandate retraining for each timestep to accommodate new labels, resulting in comparatively slower runtimes.
 6 Conclusion
-We have presented a novel simple, one-shot, and hence fast, and efficient approach to semi-supervised algorithm for graph representation learning.
-The proposed approach constructs low precision, noise-tolerant, similarity preserving randomized hyper-dimensional representations of nodes in graphs.
-We take advantage of the close connection between the Weisfeiler-Lehman test for graph isomorphism and a family of graph neural networks to design a hyper-dimensional graph learning algorithm.
-The proposed algorithm uses hyper-dimensional operators such as bundling and binding to aggregate information from the local neighborhood of each node so as to represent nodes using information that is an approximation of that used by the Weisfeiler-Lehman test.
-Results of experiments with widely used benchmark data sets show that the proposed algorithm achieves performance on semi-supervised node labeling tasks that is competitive with the state-of-the-art deep learning methods, without the need for computationally expensive optimization of model parameters and hyper-parameters.
-Consequently, the proposed method offers an especially attractive alternative with energy footprint that is substantially smaller than that of deep learning based modern graph embedding methods.
-Because the model supports one-shot learning of node embeddings, it lends itself to use in settings that call for continual learning algorithms De Lange et al.
-(2021); Wickramasinghe et al.
-(2023) wherein data samples, or classes, are presented on an ongoing basis, allowing the model to be continually updated, without the need for retraining on old data.
-Additional directions for further research include variants of the proposed algorithm that can learn graph representations that invariant or equivariant with respect to a specified set of operations on graphs Satorras et al.
-(2021); Keriven and Peyré (2019).
+We introduced HDGL, a novel and efficient algorithm for transductive learning on graphs using hyperdimensional representations.
+HDGL leverages the injectivity property of node representations of a family of Graph Neural Networks (GNNs) to map node features to the HD space and then uses HD operators such as bundling and binding to aggregate information from the local neighborhood of each node.
+The resulting latent node representations support both node classification and link prediction tasks, unlike typical deep learning methods, which often require separate models for these tasks.
+HDGL obviates the need for expensive iterative training of the sort required by deep learning methods.
+HDGL is also well-suited for class incremental learning where the model has to learn to effectively discriminate between a growing number of classes.
+Our experiments with widely used benchmark datasets show that on the node classification task, HDGL is competitive with the state-of-the-art GNN methods with respect to accuracy, at substantially reduced computational cost.
+Our experiments also show that the HD representation constructed by HDGL supports link prediction at accuracies comparable to that of DeepWalk and related methods.
+We conclude that HDGL offers a computationally efficient alternative to graph neural networks for node classification, especially in settings that call for class-incremental learning or in applications that demand high accuracy models at significantly lower computational cost and learning time than possible with the state-of-the-art graph neural network methods.

### v3

@@
-Hyperdimensional Computing for Node Classification and Link Prediction
-Abstract
-We introduce a novel method for transductive learning on graphs using hyperdimensional representations.
-The proposed approach encodes data samples using random projections into a very high-dimensional space (hyperdimensional or HD space for short).
-It obviates the need for expensive iterative training of the sort required by deep learning methods.
-Specifically, we propose a Hyperdimensional Graph Learning (HDGL) algorithm.
-HDGL leverages the injectivity property of node representations of a family of Graph Neural Networks (GNNs) to map node features to the HD space and then uses HD operators such as bundling and binding to aggregate information from the local neighborhood of each node.
-The resulting latent node representations support both node classification and link prediction tasks, unlike typical deep learning methods, which often require separate models for these tasks.
-We report results of experiments using widely used benchmark datasets which demonstrate that, on the node classification task, HDGL is competitive with the state-of-the-art GNN methods with respect to accuracy, at substantially reduced computational cost.
-Furthermore, HDGL is well-suited for class incremental learning where the model has to learn to effectively discriminate between a growing number of classes.
-Our experiments also show that the HD representation constructed by HDGL supports link prediction at accuracies comparable to that of DeepWalk and related methods, although it falls short of state-of-the-art Graph Neural Network (GNN) methods that rely on computationally expensive iterative training.
-We conclude that HDGL offers a computationally efficient alternative to graph neural networks for node classification, especially in settings that call for class-incremental learning or in applications that demand high accuracy models at significantly lower computational cost and learning time than possible with the state-of-the-art graph neural network methods.
-1 Introduction
-Node Classification and Link Prediction in Networks.
-Social networks, citation networks, etc., are naturally represented as graphs where nodes represent the entities that make up the network and edges between pairs of nodes encode relationships between entities.
-For example, in a social network, nodes represent individuals and links denote social connections; in a citation network, nodes represent articles, and (directed) links denote citations.
-Over the past decade, there has been considerable interest in graph learning problems such as node classification and link prediction [30, 29].
-Such problems are typically solved using graph embedding techniques [63, 6, 60, 55].
-1.1 Related Work
-Graph Neural Networks (GNNs), originally introduced by Sperduti and Starita [46], Frasconi et al.
-[13], Gori et al.
-[17], Scarselli et al.
-[44], which learn low-dimensional embeddings of nodes in a graph by aggregating information from local neighborhoods of the nodes, have recently emerged as the dominant approach to graph representation learning [26, 53, 18], because of their superior performance across a broad spectrum of applications [58, 62, 63].
-Like other deep learning models, GNNs are data hungry, require extensive hyperparameter tuning, and achieve their superior performance at considerable computational cost, resulting in a large carbon footprint with attendant environmental impact [10, 47, 9, 50].
-Hence, there is an urgent need for computationally more efficient, and consequently, environmentally more sustainable alternatives to deep learning models in general, with a specific emphasis on graph neural networks.
-Hyperdimensional computing [22, 23, 37, 49] offers a brain-inspired alternative to deep neural networks for artificial intelligence and machine learning applications, that maps each object using a sufficiently high-dimensional random encoding (HD encoding for short) to produce a binary or bipolar vector [22, 23].
-All computations are performed on the objects thus encoded using simple operations, e.g., element-wise additions and dot products [22, 37, 49].
-Computations on HD vectors, because they yield binary or bipolar vectors, can be realized using low-precision, fast, low-power, energy-efficient hardware [27, 28].
-Hence, there is a growing interest in the use of HD computing in artificial intelligence and machine learning [37, 28, 15, 34].
-Hyperdimensional encodings of graphs have been investigated by Poduval et al.
-[43] and Nunes et al.
-[38].
-These works focus on performing graph operations, e.g., similarity based retrieval, graph classification, etc.
-In contrast, our focus is on methods that learn node-level representations for node classification and link prediction.
-Hyperdimensional node embeddings have been explored by Kang et al.
-[24] and [31].
+Hyperdimensional Representation Learning for Node Classification and Link Prediction
+Abstract.
+We introduce Hyperdimensional Graph Learner (HDGL), a novel method for node classification and link prediction in graphs.
+HDGL maps node features into a very high-dimensional space (hyperdimensional or HD space for short) using the injectivity property of node representations in a family of Graph Neural Networks (GNNs) and then uses HD operators such as bundling and binding to aggregate information from the local neighborhood of each node yielding latent node representations that can support both node classification and link prediction tasks.
+HDGL, unlike GNNs that rely on computationally expensive iterative optimization and hyperparameter tuning, requires only a single pass through the data set.
+We report results of experiments using widely used benchmark datasets which demonstrate that, on the node classification task, HDGL achieves accuracy that is competitive with that of the state-of-the-art GNN methods at substantially reduced computational cost; and on the link prediction task, HDGL matches the performance of DeepWalk and related methods, although it falls short of computationally demanding state-of-the-art GNNs.
+Hyperdimensional Computing, Graph Neural Networks, Transductive Learning, Representation Learning
+journalyear: 2025copyright: acmlicensedconference: Proceedings of the Eighteenth ACM International Conference on Web Search and Data Mining; March 10–14, 2025; Hannover, Germanybooktitle: Proceedings of the Eighteenth ACM International Conference on Web Search and Data Mining (WSDM ’25), March 10–14, 2025, Hannover, Germanydoi: 10.1145/3701551.3703492isbn: 979-8-4007-1329-3/25/03ccs: Computing methodologies Learning latent representationsccs: Computing methodologies Semi-supervised learning settings
+1. Introduction
+Social networks, citation networks, molecular, e.g., protein-protein interaction networks, etc.
+are naturally represented as graphs where nodes represent the entities that make up the network and edges between pairs of nodes encode relationships between entities.
+For example, in a social network, nodes represent individuals and links denote social connections; in a citation network, nodes represent articles, and (directed) links denote citations; In a protein-protein interaction network, the nodes represent proteins and links represent their pairwise interactions.
+Two of the key problems presented by such data include node classification and link prediction (Li and Pi, 2020; Kumar et al., 2020).
+The state-of-the-art methods for solving such problems rely on graph representation learning methods (Zhou et al., 2022; Cai et al., 2018; Xu, 2021; Wang et al., 2022).
+The state-of-the-art performance achieved by such methods comes at the expense of high computational costs – and the large carbon footprints and the attendant adverse environmental impacts (Desislavov et al., 2023; Strubell et al., 2020; de Vries, 2023; Thompson et al., 2021) – of data-hungry deep learning methods that rely on iterative parameter optimization and hyperparameter tuning.
+Against this background, we explore a computationally efficient, one-pass learning algorithm in which the model needs to see each data sample only once, as an alternative to state-of-the-art graph neural networks for node classification and link prediction.
+Key Contributions: We introduce a Hyperdimensional Graph Learner (HDGL), a novel method for node classification and link prediction in graphs.
+HDGL maps node features into a very high-dimensional space (hyperdimensional or HD space for short) using the injectivity property of node representations in a family of Graph Neural Networks (GNNs) and then uses HD operators such as bundling and binding to aggregate information from the local neighborhood of each node yielding latent node representations that can support both node classification and link prediction tasks.
+HDGL provides a one-pass learning alternative to GNNs trained using computationally expensive iterative optimization methods.
+We report results of extensive experiments using widely used benchmark datasets which demonstrate that, on the node classification task, HDGL achieves accuracy that is competitive with that of the state-of-the-art GNN methods at substantially reduced computational cost; and on the link prediction task, HDGL matches the performance of DeepWalk and related methods, although it falls short of computationally demanding state-of-the-art GNN methods.
+We further show that HDGL, which does not require expensive iterative learning procedures, is well-suited for data/class-incremental learning, making it an attractive alternative to GNNs.
+2. Related Work
+Graph Neural Networks (GNNs), introduced by Sperduti and Starita (1997); Frasconi et al.
+(1998); Gori et al.
+(2005); Scarselli et al.
+(2009), which learn embeddings of nodes in a graph by aggregating information from local neighborhoods of the nodes, have recently emerged as the dominant approach to graph representation learning (Kipf and Welling, 2017; Veličković et al., 2018; Hamilton et al., 2017), because of their superior performance across a broad spectrum of applications (Wu et al., 2020; Zhang et al., 2018; Zhou et al., 2022).
+However, as noted above, their state-of-the-art performance comes at a high computational cost and adverse environmental impact.
+In contrast, HDGL needs only a single pass through the training data, and hence, offers a computationally efficient, low energy footprint alternative to GNN for node classification and link prediction.
+Hyperdimensional computing (Kanerva, 1988, 2009; Neubert et al., 2019; Thomas et al., 2021) offers a brain-inspired alternative to deep neural networks for artificial intelligence and machine learning applications, that maps each object using a sufficiently high-dimensional random encoding (HD encoding for short) to produce a binary or bipolar vector (Kanerva, 1988, 2009).
+Simple operations, like element-wise additions and dot products (Kanerva, 1988; Neubert et al., 2019; Thomas et al., 2021), are used to perform computations on the encoded objects.
+Computations on HD vectors, because they yield binary or bipolar vectors, can be realized using low-precision, fast, low-power, energy-efficient hardware (Kleyko et al., 2023a, b).
+Hence, there is a growing interest in the use of HD computing in machine learning (Neubert et al., 2019; Kleyko et al., 2023b; Ge and Parhi, 2020; Morris et al., 2022).
+Hyperdimensional Graph Encodings.
+Poduval et al.
+(2022) and Nunes et al.
+(2022) have investigated hyperdimensional representations of graphs for similarity-based retrieval and graph classification.
+Nunes et al.
+(2022) constructs the Graph HD vector using node features derived using PageRank (Brin and Page, 1998).
+Poduval et al.
+(2022) assign random HD-vectors to nodes and considers them as node features to build a Graph HD vector.
+Thus, both methods ignore node features, whereas our focus is on methods that learn node-level representations for node classification and link prediction.
+Hyperdimensional Node Embeddings.
+Recent work has explored hyperdimensional embeddings of nodes in a graph Kang et al.
+(2022); Li et al.
+(2023).
 Specifically, Kang et al.
-[24] proposed RelHD, a novel processing-in-memory (PIM) hardware architecture [35] based on FeFET technology [4]; and Li et al.
-[31] introduced HyperNode for learning HD embeddings of nodes.
-However, both RelHD and HyperNode suffer from significant shortcomings associated with how they construct node representations from node features (see section 3.5 for details).
-1.2 Overview and Contributions
-We introduce a novel method for transductive learning on graphs using hyperdimensional representations.
-The proposed approach encodes data samples using random projections into a very high- dimensional space (hyperdimensional or HD space for short) [22, 23, 37, 49].
-It obviates the need for expensive iterative training of the sort required by deep learning methods.
-The resulting latent node representations support both node classification and link prediction tasks, unlike typical deep learning methods, which often require separate models for these tasks.
-Contributions. The key contributions of the paper are as follows:
-  • We introduce HDGL, a method for learning HD node embeddings that match the expressive power of GNN embeddings.
-HDGL leverages the injectivity property of node representations of a family of Graph Neural Networks (GNNs) to map node features to the HD space and then uses HD operators such as bundling and binding to aggregate information from the local neighborhood of each node.
-  • We show how the HD node embeddings learned by HDGL can be used to perform both semi-supervised node labeling and link prediction tasks.
-Thus, unlike deep learning methods usually require separate models for each of these tasks, HDGL produces a single model that can support both tasks.
-  • We present results of extensive experimental comparisons of HDGL with the state-of-the-art GNN methods on several standard benchmark data sets which show that HDGL achieves performance that is competitive with that of GNN, while avoiding the need for computationally expensive training.
-  • We present results experiments which show that HDGL offers a better alternative to graph neural networks for settings that call for class-incremental learning, where the model is required to effectively discriminate between a growing number of classes that are presented incrementally over time.
-Organization.
-The rest of the paper is organized as follows: Section 3 reviews the HD computing framework, its basic operations, and their properties, and reviews the expressive power of GNN.
-Section 4 introduces HDGL, the proposed HD computing based algorithm for learning node embeddings.
-Section 5 describes results of extensive experiments that compare HDGL with the state-of-the-art GNN-based node embedding methods.
-Section 6 concludes with a brief summary of the paper and an outline of some directions for further research.
-2 Problem Statement
-Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ represents set of undirected nodes, $\mathcal{E}$ represents the set of edges $\mathcal{E}\subseteq\{(u,v)|u,v,\in\mathcal{V}\}$.
+(2022) proposed RelHD, a novel processing-in-memory (PIM) hardware architecture (Mutlu et al., 2022) based on FeFET technology (Beyer et al., 2020).
+Li et al.
+(2023) introduced the HyperNode method for learning HD embeddings of nodes.
+However, both RelHD and HyperNode encode node features as binary indicator variables before mapping them to HD space and hence are unsuited for graphs with continuous or integer-valued node attributes e.g., word counts.
+Furthermore, none of the aforementioned methods (Nunes et al., 2022; Poduval et al., 2022; Kang et al., 2022; Li et al., 2023) support link prediction on graphs.
+In contrast, HDGL offers an effective and efficient method for constructing more expressive node embeddings that can accommodate not only binary or multi-valued, but also integer-valued or continuous node features.
+The resulting HD encodings of nodes support both node classification and link prediction.
+3. Preliminaries
+We proceed to briefly summarize the key concepts needed to set the stage for introducing HDGL.
+3.1. Node Classification and Link Prediction on Graphs
+Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ represents set of nodes, $\mathcal{E}$ represents the set of undirected edges $\mathcal{E}\subseteq\{(u,v)|u,v,\in\mathcal{V}\}$.
 Let $N=|\mathcal{V}|$ denote the number of nodes in the graph.
@@
 The $k$-hop neighbors of node $v$ in graph $\mathcal{G}$ are represented using multi-set $\mathcal{N}^{k}(v)$.
-For the node label prediction problem in the transductive setting, let $y_{w}\in{1,\cdots,L}$ represent the label associated with each node $w\in\mathcal{W}$, where $\mathcal{W}\subset\mathcal{V}$.
-This implies that not all nodes in the graph are labeled.
-The goal is to learn a latent representation $\phi(v)$ for each node $v$, utilizing the structure of $\mathcal{G}$, the feature vector $\mathbf{x}_{v}$, and the labels $y_{w}$, to accurately predict the labels for a set of nodes that are unlabeled in the graph $\mathcal{G}$.In the transductive link prediction setting, the goal is to infer missing edges within the graph by leveraging the latent representations $\phi(v)$ of the nodes.
-3 Background and Preliminaries
-3.1 Properties of HD Computing
-Hyper-dimensional computing, originally introduced by Kanerva [22] is a brain-inspired approach to cognitive computing that encodes each object $x\in\mathcal{X}$ using a high dimensional mapping $\phi:\mathcal{X}\rightarrow\mathcal{H}$, where $\mathcal{H}$ is typically $\{-1,1\}^{n}$ or $\{0,1\}^{n}$ where $n$ is sufficiently large [22, 23].
-All computations are performed in $\mathcal{H}$, using simple operations, e.g., element-wise additions and dot products [22, 37].
-The mapping $\phi$ is often random, which because its elements are low precision, lends itself to fast, low-power, energy-efficient hardware realizations [27, 28].
-In this paper, we assume that =$\mathcal{H}=\{0,1\}^{n}$.
-We proceed to briefly summarize some of the key properties of HD representation and operations that can be used to compute with it.
+The node classification problem can be described as follows: Given labels $y_{w}\in\{1,\cdots,L\}$ for each node $w\in\mathcal{W}$, where $\mathcal{W}\subset\mathcal{V}$; the task is to predict the labels of unlabeled nodes, i.e., nodes in $\mathcal{V}\setminus\mathcal{W}$.
+Similarly, the link prediction problem entails predicting the missing edges/new edges in $\mathcal{G}$, e.g., predicting new links/recommendation in a social network.
+We aim to solve both problems by learning a latent representation $\phi(v)$ for each node $v$, utilizing the structure of $\mathcal{G}$, i.e., the connectivity between nodes, the feature vector $\mathbf{x}_{v}$, and the node labels $y_{w}$.
+3.2. Hyperdimensional Computing
+Hyper-dimensional computing, originally introduced by Kanerva (1988) is a brain-inspired approach to representing information that encodes each object $x\in\mathcal{X}$ using a high dimensional mapping $\phi:\mathcal{X}\rightarrow\mathcal{H}$, where $\mathcal{H}$ is typically $\{-1,1\}^{n}$ or $\{0,1\}^{n}$ where $n$ is sufficiently large (Kanerva, 1988, 2009).
+All computations are performed in $\mathcal{H}$, using simple operations, e.g., element-wise additions and dot products (Kanerva, 1988; Neubert et al., 2019).
+The mapping $\phi$ is often random, and low precision which lends itself to fast, low-power, energy-efficient hardware realizations (Kleyko et al., 2023a, b).
+In this paper, we assume =$\mathcal{H}=\{0,1\}^{n}$.
+We proceed to summarize some of the key properties of HD representation.
 Near Orthogonality of Random vectors in HD Space.
@@
 Let $\mathbf{a}\in\{0,1\}^{n}$ and $\mathbf{b}\in\{0,1\}^{n}$ be two such HD vectors where $a_{i}\sim Ber(p=0.5);\forall i\in\{1,\cdots,n\}$ and $b_{i}\sim Ber(p=0.5);\forall i\in\{1,\cdots,n\}$.
-We can then say that $\mathbf{a}$ and $\mathbf{b}$ are near orthogonal or $d_{H}[\mathbf{a},\mathbf{b}]$, the normalized hamming distance between $\mathbf{a}$ and $\mathbf{b}$, $\approx 0.5$ [22].
+We can then say that $\mathbf{a}$ and $\mathbf{b}$ are near orthogonal or $d_{H}[\mathbf{a},\mathbf{b}]$, the normalized hamming distance between $\mathbf{a}$ and $\mathbf{b}$, $\approx 0.5$ (Kanerva, 1988).
 This is also known as a blessing of HD spaces.
@@
 As noted by Neubert et al.
-[37], in the case of an $n$-dimensional binary HD space where each of the $n$-dimensional vectors is $f$-sparse, e.g., the number of nonzero entries is $fn$, the resulting HD space has capacity given by $\binom{n}{\lfloor f\cdot n\rfloor}$.
+(2019), in the case of an $n$-dimensional binary HD space where each of the $n$-dimensional vectors is $f$-sparse, e.g., the number of nonzero entries is $fn$, the resulting HD space has capacity given by $\binom{n}{\lfloor f\cdot n\rfloor}$.
 For example, when $n=1000$, and $f=0.05$, the resulting HD space can store approximately $10^{80}$ or almost as many patterns as the number of atoms in the observable universe (which is estimated to be between $10^{78}$ and $10^{82}$).
 Noise or Error Tolerance.
-Due to the distributed representation of data across numerous dimensions in HD spaces, errors in HD computing are limited to only a fraction of bits [1, 37, 49].
-Thus, HD computing offers noise tolerance without the need for expensive error correction mechanisms [37].
-3.2 HD Computing Operations
+Due to the distributed representation of data across numerous dimensions in HD spaces, errors in HD computing are limited to only a fraction of bits (Ahmad and Hawkins, 2015; Neubert et al., 2019; Thomas et al., 2021).
+Thus, HD computing offers noise tolerance without the need for expensive error correction mechanisms (Neubert et al., 2019).
+We next briefly describe some of the key operations used to compute with HD representations:-
 Bundling, denoted by $\oplus$, is a bitwise majority operation represented by :$\oplus:\{0,1\}^{n}\times\{0,1\}^{n}\rightarrow\{0,1\}^{n}$.
-Bundling two randomly sampled HD vectors results in a new vector which is similar to the input vectors.
+Bundling randomly sampled HD vectors results in a new vector which is similar to the input vectors.
 Bundling $\oplus$ is associative and commutative i.e $\mathbf{a}\oplus(\mathbf{b}\oplus\mathbf{c})=(\mathbf{a}\oplus\mathbf{b})\oplus\mathbf{c}=(\mathbf{c}\oplus\mathbf{a})\oplus\mathbf{b}$.
 Note that in case of Bundling an even number of HD vectors, there is a possibility of ties since Bundling is a bitwise majority operation.
-To break ties, a tie-breaking policy, e.g., random choice of 0 or 1 on dimensions where ties occur.
+To break ties, a tie-breaking policy needs to be used, e.g., random choice of 0 or 1 on dimensions where ties occur.
 Bundling can be thought of as the HD space analog of the mean operation euclidean spaces.
@@
 A key property of Rotation is that rotation results in dissimilar/orthogonal vectors i.e $d_{H}[\mathbf{a},\Pi(\mathbf{a})]\approx 0.5$ and it is distributive over both Bundling and Binding and like binding, rotation is also reflective i.e $d_{H}[\mathbf{a},\mathbf{b}]\approx d_{H}[\Pi(\mathbf{a}),\Pi(\mathbf{b})]$.
-Rotation was introduced by Gayler [14] in order to protect or preserve information because information can be lost/cannot be distinguished because Binding is associative and commutative i.e:- $(\mathbf{a}\otimes\mathbf{b})\otimes(\mathbf{c}\otimes\mathbf{d})=(\mathbf{a}\otimes\mathbf{c})\otimes(\mathbf{b}\otimes\mathbf{d})$.
+Rotation was introduced by Gayler (1998) in order to protect or preserve information because information can be lost/cannot be distinguished because Binding is associative and commutative i.e:- $(\mathbf{a}\otimes\mathbf{b})\otimes(\mathbf{c}\otimes\mathbf{d})=(\mathbf{a}\otimes\mathbf{c})\otimes(\mathbf{b}\otimes\mathbf{d})$.
 In order to make the make a distinction between $(\mathbf{a}\otimes\mathbf{b})\otimes(\mathbf{c}\otimes\mathbf{d})$ and $(\mathbf{a}\otimes\mathbf{c})\otimes(\mathbf{b}\otimes\mathbf{d})$, rotation is used :- $(\mathbf{a}\otimes\mathbf{b})\otimes\Pi(\mathbf{c}\otimes\mathbf{d})$.
@@
 Specifically, the Bundle operation can be represented as signed addition in the bipolar space, while the binding operation aligns with multiplication in the bipolar space.
-3.3 GNN-based Node Embdeddings
+3.3. Graph Neural Networks
 […]
-3.4 Representational Power of GNN
-Several authors have investigated the representational power of GNN, including their ability to distinguish between different graph structures [59], e.g., using the Weisfeiler-Lehman (WL) graph isomorphism test [56, 19] , or to learn a set of functions on graphs [44], or the equivalence of the preceding two notions of expressivity[8].
-The WL test iteratively updates a node’s feature vector by aggregating the feature vectors of its network neighbors [56, 2].
-The injectivity of the aggregation function used ensures that different neighborhoods of the nodes map to different feature vectors [19].
-Alternatively, if the set of feature vectors of a node are represented as a multiset, the aggregation function must be able to map different multisets into distinct encodings.
-As shown by [59, 33], GNN are at most as powerful as the WL test in distinguishing graph structures if the mapping $\mathcal{A}:\mathcal{G}\rightarrow\mathbb{R}^{d}$ produced by the iterative aggregation scheme: $\mathbf{h}^{(k)}_{v}=g(\mathbf{h}^{(k-1)}_{v},f(\{\mathbf{h}^{(k-1)}_{j};j\in\mathcal{N}^{1}(v)\}))$ where $g(\cdot)$ and $f(\cdot)$ and $\mathcal{A}$’s graph-level readout are injective.
-GNN generalize the WL test in that, in addition to discriminating different graph structures as they learn to map similar graph structures to similar representations [59].
-3.5 Limitations of Previous HD Graph Methods
-The HD Computing based graph-level learning methods introduced by Nunes et al.
-[38] and Poduval et al.
-[43] are designed for similarity-based graph retrieval and graph classification, as opposed to node classification or link prediction.
-Nunes et al.
-[38] constructs the Graph HD vector using node features derived using PageRank [5].
-Poduval et al.
-[43] assign random HD-vectors to nodes and considers them as node features to build a Graph HD vector.
-Thus, both methods ignore node features present in the graph.
-RelHD [24] constructs node representations for node classification tasks and uses the resulting representation to design a novel hardware chip design with a scalable processing in-memory (PIM) architecture based FeFET.
-However, RelHD has a significant limitation with respect how it encodes the node features into HD vectors.
-Specifically, it treats node features as dimension-wise indicator variables and computes the HD vector representation of each node by first generating $d$ random basis vector.
-Let $\Upsilon_{p}$ be such a random basis Hyper vector.
-Random basis vectors $\Upsilon_{p}$ are bundled/accumulated if feature at index $p$ is present i.e HD vector for a node $u$ is computed using basis vectors and raw features $\mathbf{x}_{u}=[x_{u1},\ldots,x_{ud}]$ using the following formula:- $\oplus_{p\in\{1,\ldots,d\}}(\mathbbm{1}(x_{up}\neq 0)\cdot\Upsilon_{p})$.
-Consequently, this method is not suitable for graphs with continuous or integer-valued node attributes in continuous form or even simple bag-of-words counts.
-Note that because HyperNode [31] also utilizes the same mechanism for constructing HD vector encodings of the nodes, it suffers from the same limitations as RelHD.
-Furthermore, none of the aforementioned methods [38, 43, 24, 31] support link prediction on graphs.
-In contrast, HDGL offers an effective and efficient method for constructing more expressive node embeddings that can accommodate not only binary or multi-valued, but also integer-valued or continuous node features.
-The resulting HD encodings of nodes support both node classification and link prediction.
-4 Hyper-Dimensional Graph Learner
-In essence, our design of HDGL aims to achieve a high level of expressive power similar to graph neural networks (GNN).
-We utilize HD computing operations on node representations, taking advantage of HD operations—Bundling ($\oplus$), Binding ($\otimes$), and Rotation ($\Pi$) —and their properties.
-This ensures that HDGL, akin to GNN, not only aspires to build expressive power but also learns to map similar graphs to similar representations.
-The proposed HD learner carries out 3 major steps: (i).
-Map Node features $\mathbf{x}_{v}\in\mathbb{R}^{d}$ to a intermediate HD representation $\mathbf{r}_{v}\in\{0,1\}^{\beta}$ $\forall v\in\mathcal{V}$.
+Representational Power of GNN.
+The representational power of GNNs can be characterized in terms of their ability to distinguish between different graph structures (Xu et al., 2019).
+The greater the representational power, the more minute the differences between graphs that can be detected by the GNN.
+GNN have been shown to be at most as powerful as the Weisfeiler-Lehman (WL) graph isomorphism test, which iteratively updates node features by aggregating neighbors’ features using an injective aggregation function (Weisfeiler and Leman, 1968; Huang and Villar, 2021).
+The injectivity of the aggregation function ensures that the differences in node neighborhoods result in differences in the resulting aggregations (Huang and Villar, 2021).
+Consequently, GNN are at most as powerful as the WL test in distinguishing graph structures if the mapping $\mathcal{A}:\mathcal{G}\rightarrow\mathbb{R}^{d}$ produced by the iterative aggregation scheme: $\mathbf{h}^{(k)}_{v}=g(\mathbf{h}^{(k-1)}_{v},f(\{\mathbf{h}^{(k-1)}_{j};j\in\mathcal{N}^{1}(v)\}))$ where $g(\cdot)$ and $f(\cdot)$ and $\mathcal{A}$’s graph-level readout are injective.
+4. Hyper-Dimensional Graph Learner
+HDGL constructs a HD representation of graphs that matches the expressive power of GNN.
+HDGL exploits the properties of HD operations, namely, Bundling ($\oplus$), Binding ($\otimes$), and Rotation ($\Pi$) to ensure that graphs with similar node features and topologies map to similar HD representations.
+Specifically, HDGL: (i).
+Maps node features $\mathbf{x}_{v}\in\mathbb{R}^{d}$ to a intermediate HD representation $\mathbf{r}_{v}\in\{0,1\}^{\beta}$ $\forall v\in\mathcal{V}$.
 (ii).
-Construct a latent HD representation $\mathbf{z}_{v}\in\{0,1\}^{\beta}$ using $\mathbf{r}_{v}$ and $\mathcal{N}^{k}(v)$.
+Constructs a latent HD representation $\mathbf{z}_{v}\in\{0,1\}^{\beta}$ using $\mathbf{r}_{v}$ and $\mathcal{N}^{k}(v)$.
 (iii).
-Use $\mathbf{z}_{v}$ to predict $y_{v}$ or predict links/edges in the graph.
+Uses $\mathbf{z}_{v}$ to predict node labels and links in the graph.
 We proceed to describe each of these steps in detail.
-Figure 1: An example mapping node features to a binary space using Random Hyperplanes.
-A point receives a value of 1 when on the left side of a hyperplane and 0 when on the right side.
-Utilizing multiple hyperplanes results in a binary sketch for each point.
-4.1 Mapping Node Features to HD-space
+Figure 1. Mapping node features to a binary space using Random Hyperplanes: points are assigned 1 or 0 based on their position relative to each hyperplane.
+[…]
+4.1. Mapping Node Features to HD-space
 The first crucial step is to map node features from $\mathbb{R}^{d}$ space to $\{0,1\}^{\beta}$ HD space such that $\beta>>d$.
-We do this using random hyperplane tessellations [12], a type of Locality Sensitive Hashing [21, 7].
-The key idea is to generate $\beta$ random hyperplanes in $\mathbb{R}^{d}$ and set the corresponding co-ordinate of the HD representation of a node based on whether the its feature vector in $\mathbb{R}^{d}$ lies on the positive or the negative side of the hyperplane.
+We do this using random hyperplane tessellations (Dirksen et al., 2022), a type of Locality Sensitive Hashing (Indyk and Motwani, 1998; Charikar, 2002).
+The key idea is to generate $\beta$ random hyperplanes in $\mathbb{R}^{d}$ and set the corresponding co-ordinate of the HD representation of a node based on whether its feature vector in $\mathbb{R}^{d}$ lies on the positive or the negative side of the hyperplane.
 With $\beta$ such hyper-planes in hand, we can generate an $n$ dimensional binary sketch for each node.
-More precisely, let $\mathbf{A}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{\beta}$ be a matrix with each row $\mathbf{a}_{1}^{T},\cdots\mathbf{a}_{\beta}^{T}$ drawn from $\mathcal{N}(0,I_{d})$ and $\gamma$ be uniformly distributed on $[-\lambda,\lambda]$.
+More precisely, let $\mathbf{Q}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{\beta}$ be a matrix with each row $\mathbf{q}_{1}^{T},\cdots\mathbf{q}_{\beta}^{T}$ drawn from $\mathcal{N}(0,I_{d})$ and $\gamma$ be uniformly distributed on $[-\lambda,\lambda]$.
 For conciseness, we represent $\beta$ samples of $\gamma$ as $\Gamma\sim[-\lambda,\lambda]^{\beta}$.
-Using $\mathbf{A}$ and $\Gamma$, a randomized binary sketch $\mathbf{r}_{v}$ for a node $v\in\mathcal{V}$ is given by $\mathbf{r}_{v}=\text{sign}(\mathbf{A}\mathbf{x}_{v}+\Gamma)$.
-The primary purpose of adding the $\Gamma$ offset is to ensure that the not all sampled hyperplanes pass through the origin.
-When dealing with evenly spread points, typically in the context of real-valued vectors, setting $\Gamma$ to zero is often appropriate.
+Using $\mathbf{Q}$ and $\Gamma$, a randomized binary sketch $\mathbf{r}_{v}$ for a node $v\in\mathcal{V}$ is given by $\mathbf{r}_{v}=\text{sign}(\mathbf{Q}\mathbf{x}_{v}+\Gamma)$.
+The offset $\Gamma$ when it is non-zero ensures that the sampled hyperplane does not pass through the origin.
 Note that the resulting binary sketch of each node is in fact a $\beta$-dimensional binary HD vector.
 Nodes with similar features will have similar HD representations; and nodes with dissimilar features will have dissimilar HD representations.
-4.2 Neighborhood Aggregation Scheme
-Under the simplifying assumption that nodes that have identical multi-sets of features over their $k$-hop neighborhoods are likely to also have matching local topologies, we approximate the injectivity of the embedding $\mathbf{z}_{v}$ which maps each node to its HD representation by aggregating features of nodes over a $k$-hop neighborhood of $v$ as follows: $\mathbf{z}_{v}=\phi\Big{(}\mathbf{r}_{v},\varphi_{1}(\{\mathbf{r}_{j};j\in\mathcal{N}^{1}(v)\}),\ldots,\varphi_{K}(\{\mathbf{r}_{m};m\in\mathcal{N}^{K}(v)\})\Big{)}$ where $\mathbf{z}_{v}$ denotes the HD embedding of node $v$ and $\phi(\cdot)$ and $\varphi_{k}(\cdot);k\in\{1,\cdots,K\}$ are injective functions.
-Bundling for generalizing across similar nodes.
-We propose that for a node $v\in\mathcal{V}$; the functions $\varphi_{k}(\cdot)$ should be the Bundling operation $\oplus$ since, due to its permutation invariance, allows us to represent unordered sets (or multi-sets).
-Hence, we propose to use Bundling to realize $\varphi_{k}$.
-More precisely, $\varphi_{k}\Big{(}\{\mathbf{r}_{j};j\in\mathcal{N}^{k}(v)\}\Big{)}=\mathop{\oplus}_{j\in\mathcal{N}^{k}(v)}\mathbf{r}_{j}$ where $\mathop{\oplus}_{j\in\mathcal{N}^{k}(v)}\mathbf{r}_{j}$ denotes binding the HD representations of all of the $k$-hop neighbors of $v$.
-This operation is analogous to finding a high-dimensional vector that is the most similar to all of the input HD-vectors, akin to how we find the mean in Euclidean spaces and like the mean, it is not injective.
-However, as noted by [59], GNNs that use the mean to aggregate features over node neighborhoods [26, 18] perform quite well in practice because the mean operation becomes increasingly injective when graphs consist of diverse node features, which is often the case in real-world graphs.
-It is worth noting that the Binding operation $\otimes$ would be inappropriate in this context because binding of two similar vectors results in dissimilar vectors.
-Thus, if Binding were used to realize $\varphi_{k}$, similar yet distinct local neighborhoods structures would yield different HD representations, adversely impacting HDGL’s ability to generalize across such graph structures.
-To complicate matters, Binding is also invertible i.e $\mathbf{a}\otimes\mathbf{b}\otimes\mathbf{b}=\mathbf{a}$; and hence, it could discard information that is useful for downstream task e.g., node label prediction.
-Binding for linking node with its neighborhood.
-Having incorporated the Bundling operation in our framework by assigning it to the function $\varphi_{k}(\cdot)$, the remaining task involves identifying a suitable operation for the function $\phi(\cdot)$, for coupling the representation of a node obtained using bundling (see above) with the representations of nodes in its neighborhood.
-We propose to use Binding the operation $\otimes$ to realize the function $\phi(\cdot)$ because it offers a means to balance the discriminative power of the resulting model against its generalization ability.
-We illustrate the preceding claim with an example.
+4.2. Neighborhood Aggregation Scheme
+To simplify the mapping of nodes to their HD encodings, we assume that nodes that have identical multi-sets of features over their $k$-hop neighborhoods are likely to also have identical local topologies.
+Under this assumption, we can approximate injectivity of the embedding $\mathbf{z}_{v}$ of each node $v$ to its HD representation by aggregating features of nodes over a $k$-hop neighborhood of $v$ as follows: $\mathbf{z}_{v}=\phi\Big{(}\mathbf{r}_{v},\varphi_{1}(\{\mathbf{r}_{j};j\in\mathcal{N}^{1}(v)\}),\ldots,\varphi_{K}(\{\mathbf{r}_{m};m\in\mathcal{N}^{K}(v)\})\Big{)}$ where $\mathbf{z}_{v}$ denotes the HD embedding of node $v$ and $\phi(\cdot)$ and $\varphi_{k}(\cdot);k\in\{1,\cdots,K\}$ are injective functions.
+Ensuring that similar nodes have similar HD representations.
+Suppose for a moment that the Binding operation $\otimes$ is used to realize $\varphi_{k}(\cdot)$.
+Because binding of two similar vectors results in dissimilar vectors, it would adversely impact the utility of the HD representation for predicting node labels based on node features and local node topologies.
+In contrast, we use Bundling operation because it is permutation-invariant and is well-suited for implementing the functions $\varphi_{k}(\cdot)$ under the assumption that nodes that have identical multi-sets of features over their $k$-hop neighborhoods are likely to also have identical local topologies.
+More precisely, $\varphi_{k}\Big{(}\{\mathbf{r}_{j};j\in\mathcal{N}^{k}(v)\}\Big{)}=\mathop{\oplus}_{j\in\mathcal{N}^{k}(v)}\mathbf{r}_{j}$ where $\mathop{\oplus}_{j\in\mathcal{N}^{k}(v)}\mathbf{r}_{j}$ denotes bundling the HD representations of all of the $k$-hop neighbors of $v$.
+This operation is analogous to finding a high-dimensional vector that is most similar to all of the input high-dimensional vectors, similar to finding the mean in Euclidean spaces and like the mean, it is not necessarily injective.
+However, as noted by (Xu et al., 2019), GNNs that use the mean to aggregate features over node neighborhoods (Kipf and Welling, 2017; Hamilton et al., 2017) perform quite well in practice because the mean operation becomes increasingly injective when graphs consist of diverse node features, which is typically the case in real-world graphs.
+Binding for linking node with its neighbors.
+Having chosen Bundling to realize $\varphi_{k}(\cdot)$, we proceed to consider how to realize $\phi(\cdot)$, which has to couple the representation of a node obtained using bundling (see above) with the representations of its neighbors.
+As we will see shortly, if we choose to realize $\phi(\cdot)$ using the Binding operation $\otimes$, we obtain a representation that strikes a reasonable balance between discriminating between dissimilar graphs and generalizing over similar graphs.
 Let $v\in\mathcal{V}$ and $w\in\mathcal{V}$ be two nodes in a graph.
@@
 Considering only 1-hop and 2-hop neighbors of $v$, the latent representation of $v$ and $w$ will be
-$\begin{split}\mathbf{z}_{v}&=\mathbf{r}_{v}\otimes\Big{(}\mathop{\oplus}_{j\in\mathcal{N}^{1}(v)}\mathbf{r}_{j}\Big{)}\otimes\Big{(}\mathop{\oplus}_{m\in\mathcal{N}^{2}(v)}\mathbf{r}_{m}\Big{)}\\ \mathbf{z}_{w}&=\mathbf{r}_{w}\otimes\Big{(}\mathop{\oplus}_{j\in\mathcal{N}^{1}(w)}\mathbf{r}_{j}\Big{)}\otimes\Big{(}\mathop{\oplus}_{m\in\mathcal{N}^{2}(w)}\mathbf{r}_{m}\Big{)}\end{split}$ (1)
+(1) $\begin{split}\mathbf{z}_{v}&=\mathbf{r}_{v}\otimes\Big{(}\mathop{\oplus}_{j\in\mathcal{N}^{1}(v)}\mathbf{r}_{j}\Big{)}\otimes\Big{(}\mathop{\oplus}_{m\in\mathcal{N}^{2}(v)}\mathbf{r}_{m}\Big{)}\\ \mathbf{z}_{w}&=\mathbf{r}_{w}\otimes\Big{(}\mathop{\oplus}_{j\in\mathcal{N}^{1}(w)}\mathbf{r}_{j}\Big{)}\otimes\Big{(}\mathop{\oplus}_{m\in\mathcal{N}^{2}(w)}\mathbf{r}_{m}\Big{)}\end{split}$
 Suppose we have
@@
 It is easy to see that if all the terms in $\mathbf{z}_{v}$ and $\mathbf{z}_{w}$ are similar, i.e., $v$ and $w$ have similar node features, and similar 1-hop and 2-hop neighborhoods, then $\mathbf{z}_{v}\approx\mathbf{z}_{w}$ as desired.
-On the other hand, if the encodings of the features of the 2-hop neighbors of both $v$ and $w$ differ, it results in the dissimilarity between $\mathbf{z}_{v}$ and $\mathbf{z}_{w}$ a desired outcome.
+On the other hand, if the encodings of the features of the 2-hop neighbors of both $v$ and $w$ differ, $\mathbf{z}_{v}$ and $\mathbf{z}_{w}$ will be dissimilar as desired.
 When not only the 2-hop neighbors but also the 1-hop neighbors of both $v$ and $w$ are dissimilar, the dissimilarity between $\mathbf{z}_{v}$ and $\mathbf{z}_{w}$ becomes even greater.
 The key property that makes binding attractive for realizing the function $\phi(\cdot)$ is its reflectivity, i.e., the distance between two HD vectors remains unchanged when both of them are bound to the same third vector.
-Conversely, it is easy to see that Bundling is not a good choice for realizing $\phi(\cdot)$.
-Bundling of the node vector $\mathbf{r}_{v}$ with neighborhood information vector $\oplus_{j\in\mathcal{N}^{k}(v)}\mathbf{r}_{j}$, a vector that has already been bundled with $k$-hop neighbors, will result in a vector that predominantly represents information from a node’s neighborhood, while leaving out essential information about the itself.
-This is due to the fact that Bundling is a bitwise majority operation; and hence, if bundling is used to realize $\phi(\cdot)$, the information from a node’s neighborhood suppresses the information about the node itself.
-Table 1: Dataset Summary.
-Dataset Nodes Edges Features Feature Type Classes Label Rate Edge Density
-CORA 2708 5429 1433 $\{0,1\}$ 7 0.052 0.0004
-CiteSeer 3327 4732 3703 $\{0,1\}$ 6 0.036 0.0004
-PubMed 19717 44338 500 $(0,1)$ 3 0.003 0.0001
-BlogCatalog 5196 171743 8189 $\mathbb{Z}^{+}$ 6 0.023 0.0127
-Coauthor CS 18333 81894 6805 $\mathbb{Z}^{+}$ 15 0.016 0.0001
-Coauthor Physics 34493 247962 8415 $\mathbb{Z}^{+}$ 5 0.002 0.0001
-DBLP 17725 105781 768 $\mathbb{R}$ 4 0.004 0.0006
+On the other hand, it is easy to see that Bundling is not a good choice for realizing $\phi(\cdot)$.
+Because bundling is a bitwise majority operation, bundling the node vector $\mathbf{r}_{v}$ with neighborhood information vector $\oplus_{j\in\mathcal{N}^{k}(v)}\mathbf{r}_{j}$; results in a vector where the information from a node’s neighborhood suppresses the information about the node’s features, which is an undesirable property.
 Improving injectivity via Rotation.
-Because Binding is associative and commutative, it can adversly impact the injectivity and consequently, therepresentational power of our HD model.
+Because Binding is associative and commutative, it can adversely impact the injectivity of $\varphi_{k}(\cdot)$ and hence the representational power of our HD model.
 For example, suppose $v\in\mathcal{V}$ and $w\in\mathcal{V}$ are two nodes in a graph with encoding obtained using HD mapping that uses Bundling to realize $\varphi_{k}$ and binding to realize $\phi(\cdot)$ and $\mathbf{z}_{v}$ and $\mathbf{z}_{w}$ are computed as per Eq.
@@
 Now, assuming $\mathcal{N}^{1}(v)\neq\mathcal{N}^{2}(v)$, the representation of $w$ can be rewritten as $\mathbf{z}_{w}=\mathbf{r}_{v}\otimes\Big{(}\mathop{\oplus}_{j\in\mathcal{N}^{2}(v)}\mathbf{r}_{j}\Big{)}\otimes\Big{(}\mathop{\oplus}_{m\in\mathcal{N}^{1}(v)}\mathbf{r}_{m}\Big{)}$.
-Since Binding is both associative and commutative, the latent representations $\mathbf{z}_{v}$ and $\mathbf{z}_{w}$ become equal, despite their 1-hop and 2-hop neighbors being different.
-This poses a problem is terms of discriminative power of the model, leading to a lack of expressive power in our HD-algorithm.
-To address this issue, we utilize the Rotation $\Pi$ operator as it protects/preserves information.
-To be more specific, with the incorporation of the rotation/permutation operator, the latent representation for nodes $v\in\mathcal{V}$ and $w\in\mathcal{V}$ changes to
+Since, Binding is both associative and commutative, the latent representations $\mathbf{z}_{v}$ and $\mathbf{z}_{w}$ become equal, despite their 1-hop and 2-hop neighbors being different.
+This adversely impacts the representation’s ability to discriminate between dissimilar graphs.
+To address this issue, we utilize the Rotation $\Pi$ operator.
+Now, the latent representation for nodes $v\in\mathcal{V}$ and $w\in\mathcal{V}$ changes:
 $\begin{split}\mathbf{z}_{v}=\mathbf{r}_{v}\otimes\Pi\Big{(}\mathop{\oplus}_{j\in\mathcal{N}^{1}(v)}\mathbf{r}_{j}\Big{)}\otimes\Pi\;\Pi\Big{(}\mathop{\oplus}_{m\in\mathcal{N}^{2}(v)}\mathbf{r}_{m}\Big{)}\\ \mathbf{z}_{w}=\mathbf{r}_{v}\otimes\Pi\Big{(}\mathop{\oplus}_{j\in\mathcal{N}^{2}(v)}\mathbf{r}_{j}\Big{)}\otimes\Pi\;\Pi\Big{(}\mathop{\oplus}_{m\in\mathcal{N}^{1}(v)}\mathbf{r}_{m}\Big{)}\end{split}$
 where $\Pi(\cdot)$ and $\Pi\;\Pi(\cdot)$ denote single and double rotation operations respectively.
-The preceding equation ensures that $\mathbf{z}_{v}$ and $\mathbf{z}_{w}$ will be dissimilar since rotation of a HD vector yields a HD vector that is dissimilar/orthogonal to the original.
+Because the rotation of HD vector yields a HD vector that is dissimilar or orthogonal to the original, now $\mathbf{z}_{v}$ and $\mathbf{z}_{w}$ will be dissimilar.
 That is,
@@
 Hence, considering only 1-hop and 2-hop neighbors, the proposed aggregation scheme to be used by HDGL for computing the HD representation of a node $v\in\mathcal{V}$ in the graph is given by
-$\begin{split}\mathbf{z}_{v}=\mathbf{r}_{v}\otimes\Pi\Big{(}\mathop{\oplus}_{j\in\mathcal{N}^{1}(v)}\mathbf{r}_{j}\Big{)}\otimes\Pi\;\Pi\Big{(}\mathop{\oplus}_{m\in\mathcal{N}^{2}(v)}\mathbf{r}_{m}\Big{)}\\ \end{split}$ (2)
-4.3 Node Label Prediction using HDGL
-With the procedure by which HDGL constructs the HD representation for each of the nodes in a graph, we turn to the task of using the resulting HD node representation to predict the node labels using $\mathbf{z}_{v}$.
-Here, we follow the standard approach to predicting class labels of HD encoded objects [15].
+(2) $\begin{split}\mathbf{z}_{v}=\mathbf{r}_{v}\otimes\Pi\Big{(}\mathop{\oplus}_{j\in\mathcal{N}^{1}(v)}\mathbf{r}_{j}\Big{)}\otimes\Pi\;\Pi\Big{(}\mathop{\oplus}_{m\in\mathcal{N}^{2}(v)}\mathbf{r}_{m}\Big{)}\\ \end{split}$
+Input: $\mathcal{G}=(\mathcal{V},\mathcal{E})$, Node Features $\mathbf{x}_{v}$, Train Nodes $\mathcal{T}_{\ell}\subset\mathcal{V}$ for each class $\forall\ell\in\{1,\cdots,L\}$ and Unlabeled Test Nodes $\mathcal{V}_{\text{test}}$
+Sample $\mathbf{Q}:[\mathbf{q}_{1}^{T}\sim\mathcal{N}(0,I_{d}),\cdots,\mathbf{q}_{\beta}^{T}\sim\mathcal{N}(0,I_{d})]$;
+Sample $\Gamma\sim[-\lambda,\lambda]^{\beta}$;
+for $v\in\mathcal{V}$ do
+       $\mathbf{r}_{v}\leftarrow\text{sign}(\mathbf{Q}\mathbf{x}_{v}+\Gamma)$;
+for $v\in\mathcal{V}$ do
+       $\mathbf{z}_{v}\leftarrow\mathbf{r}_{v}\otimes\Pi\Big{(}\mathop{\oplus}_{j\in\mathcal{N}^{1}(v)}\mathbf{r}_{j}\Big{)}\otimes\Pi\;\Pi\Big{(}\mathop{\oplus}_{m\in\mathcal{N}^{2}(v)}\mathbf{r}_{m}\Big{)}$;
+if Node Label Prediction Task then
+      for $\ell\in\{1,\cdots,L\}$ do
+             $\mathbf{c}_{\ell}\leftarrow\mathop{\oplus}_{i\in\mathcal{T}_{\ell}}\mathbf{z}_{i};\quad$Computing Label Hyper Vectors;
+      for $u\in\mathcal{V}_{\text{test}}$ do
+             $\hat{y}_{u}\leftarrow\mathop{\operatorname{argmin}}_{\ell\in\{1,\cdots,L\}}d_{H}[\mathbf{z}_{u},\mathbf{c}_{\ell}];\>$Test-set Inference;
+if Link Prediction Task then
+       $\mathbf{e}^{+}\leftarrow\mathop{\oplus}_{(u,v)\in\mathcal{E}}(\mathbf{z}_{u}\otimes\mathbf{z}_{v})$ and $\mathbf{e}^{-}\leftarrow\mathop{\oplus}_{(u,v)\notin\mathcal{E}}(\mathbf{z}_{u}\otimes\mathbf{z}_{v})$;
+       Compute $\mathbf{Z}^{+}$ and $\mathbf{Z}^{-}$ using Eqn. (3);
+       $\mathbf{D}^{+}\leftarrow dist(\mathbf{Z}^{+},\mathbf{Z})$ and $\mathbf{D}^{-}\leftarrow dist(\mathbf{Z}^{-},\mathbf{Z})$;
+       for $i=1$ to $N$  do
+             for $j=1$ to $N$ do
+                   if $D^{+}_{ij}<D^{-}_{ij}$ then
+                         $\hat{A}_{ij}\leftarrow\sigma\big{(}(1-D^{+}_{ij})+D^{-}_{ij}\big{)}$;
+                  else
+                         $\hat{A}_{ij}\leftarrow\sigma\big{(}D^{+}_{ij}-(1-D^{-}_{ij})\big{)}$;
+Algorithm 1 Pseudo-code for HDGL
+4.3. Node Label Prediction using HDGL
+With the procedure for constructing a HD representation of nodes and their local topologies in place, we turn to the task of using the resulting HD node representation to predict the node labels using $\mathbf{z}_{v}$.
+Here, we follow the standard approach to predicting class labels of HD encoded objects (Ge and Parhi, 2020).
 Specifically, during the learning phase, we Bundle the collection of HD encodings of nodes belonging to each class yielding a HD Class Hyper-Vector for each of the classes.
@@
 More precisely let $\forall\ell\in\{1,\cdots,L\}$, $\mathcal{T}_{\ell}\subset\mathcal{V}$ be the subset of nodes that make up the training set with label $\ell$ and $\mathcal{V}_{\text{test}}\subset\mathcal{V}$ be the subset of unlabeled nodes in the graph.
-The class HD are constructed for each class as follows: $\mathbf{c}_{\ell}=\mathop{\oplus}_{i\in\mathcal{T}_{\ell}}\mathbf{z}_{i}\quad;\;\forall\ell\in\{1,\cdots,L\}$, resulting in a set of class HD vectors $\{\mathbf{c}_{1},\cdots,\mathbf{c}_{L}\}$.
+The class HD vectors are constructed for each class as follows: $\mathbf{c}_{\ell}=\mathop{\oplus}_{i\in\mathcal{T}_{\ell}}\mathbf{z}_{i}\quad;\;\forall\ell\in\{1,\cdots,L\}$, resulting in a set of class HD vectors $\{\mathbf{c}_{1},\cdots,\mathbf{c}_{L}\}$.
 Let $u\in\mathcal{V}_{\text{test}}$ be a node for which the class label is to be predicted.
 We predict the class label for $u$ using its HD representation and the HD representations of each of the classes as follows: $\hat{y}_{u}=\mathop{\operatorname{argmin}}_{\ell\in\{1,\cdots,L\}}d_{H}[\mathbf{z}_{u},\mathbf{c}_{\ell}]$.
-4.4 Link Prediction Using HDGL
-Section 4.1 and 4.2 outline a technique for constructing node representations optimized for node classification.
-Here we show that the same node representation can be used to support the transductive link prediction task.
-Kanerva [22, 23] describes a method for encoding entries of a dataset.
-Let $\mathbf{a}_{1}$, $\mathbf{a}_{2}$, $\mathbf{a}_{3}$ be some randomly sampled HD vectors representing data entry attributes (eg.
-weight, height and gender).
-Let $\mathbf{\alpha}$, $\mathbf{\beta}$ and $\mathbf{\gamma}$ be values for the corresponding attributes represented in HD format.
-Let $\mathbf{r}$ be a HD vector such that: $\mathbf{r}=(\mathbf{a}_{1}\otimes\mathbf{\alpha})\oplus(\mathbf{a}_{2}\otimes\mathbf{\beta})\oplus(\mathbf{a}_{3}\otimes\mathbf{\gamma}).$ From the properties of HD-computing [22], it is possible to retrieve values for a corresponding attribute using the Binding $\otimes$ operator.
-Specifically, $\mathbf{\alpha}\approx\mathbf{r}\otimes\mathbf{a}_{1};\quad\mathbf{\beta}\approx\mathbf{r}\otimes\mathbf{a}_{2};\quad\mathbf{\gamma}\approx\mathbf{r}\otimes\mathbf{a}_{3};$
-We apply the same concept to link prediction.
-Let $\mathbf{e}^{+}$ be a hypervector storing information about all or some of the edges present in the graph, and let $\mathbf{e}^{-}$ be a hypervector representing information about some of the non-edges in the graph.
-Formally: $\mathbf{e}^{+}=\mathop{\oplus}_{(u,v)\in\mathcal{E}}(\mathbf{z}_{u}\otimes\mathbf{z}_{v});\quad\mathbf{e}^{-}=\mathop{\oplus}_{(u,v)\notin\mathcal{E}}(\mathbf{z}_{u}\otimes\mathbf{z}_{v}).$ Let $\mathbf{Z}$ have $N$ rows, storing latent representations of all the nodes in the graph.
-That is, a row in $\mathbf{Z}$ is represented by $\mathbf{z}_{u}$, where $u\in\mathcal{V}$.
-Now that we have encoded edges and non-edges in the graph, the goal is to find new or missing links in the graph.
-This can also be thought of as reconstructing/generating the adjacency matrix.
-To find possible new links we Bind $\otimes$ the edge information hypervector $\mathbf{e}^{+}$ and the latent representation of all the nodes in the graph.
-This results in a new matrix $\mathbf{Z}^{+}$ where each row is $\mathbf{z}_{u}\otimes\mathbf{e}^{+}$ where $u\in\mathcal{V}$.
-$\mathbf{Z}^{+}$ represents the possible node representations of the target nodes in a link, with the corresponding source node representation being $\mathbf{Z}$.
-We follow the same procedure for the non-edge hypervector, resulting in $\mathbf{Z}^{-}$.
+4.4. Link Prediction Using HDGL
+Our approach takes advantage of a technique introduced by Kanerva (1988, 2009)’s for representing and retrieving data using HD representation.
+Lets consider a graph with 7 nodes and there are 3 edges in the graph betweem node $(1,2)$, $(3,6)$ and $(5,7)$.
+By binding latent representation between pairs of nodes which have edges between them and bundling the resulting vectors, we obtain a sum-vector that encapsulates the edge information of the graph: $\mathbf{e}^{+}=(\mathbf{z}_{1}\otimes\mathbf{z}_{2})\oplus(\mathbf{z}_{3}\otimes\mathbf{z}_{6})\oplus(\mathbf{z}_{5}\otimes\mathbf{z}_{7})$.
+Edge information can be recovered solely using hypervectors $\mathbf{e}^{+}$ and $\mathbf{z}_{i};\forall i\in\mathcal{V}$.
+For eg.
+$\mathbf{z}_{2}\approx\mathbf{e}^{+}\otimes\mathbf{z}_{1}$ and $\mathbf{z}_{7}\approx\mathbf{e}^{+}\otimes\mathbf{z}_{5}$.
+For the transductive link prediction problem, we leverage on the assumption that the underlying semantics governing existing links, which are based on latent representations of nodes in an edge, will also apply to new edges.
+For instance, if nodes $(1,2)$ are connected and we are predicting edges involving node $4$ and if $\mathbf{z}_{1}\approx\mathbf{z}_{4}$, then we would expect an edge $(4,2)$ to be predicted.
+Note that it’s also crucial to accurately predict non-edges to minimize false positives.
+Therefore, we also make use of hypervector $\mathbf{e}^{-}$ to represent information about edges that are absent in the graph.
+Following this overview, we can now delve into a detailed explanation of our model.
+Formally: $\mathbf{e}^{+}=\mathop{\oplus}_{(u,v)\in\mathcal{E}}(\mathbf{z}_{u}\otimes\mathbf{z}_{v});\quad\mathbf{e}^{-}=\mathop{\oplus}_{(u,v)\notin\mathcal{E}}(\mathbf{z}_{u}\otimes\mathbf{z}_{v}).$ Given this encoding of edges that are present and the edges that are absent in the graph, we can proceed to predict links as follows:
+To find putative new links we Bind $\otimes$ the edge information hypervector $\mathbf{e}^{+}$ and the latent representation of all the nodes in the graph.
+This results in a new matrix $\mathbf{Z}^{+}$ where the $i^{th}$ row is $\mathbf{z}_{i}\otimes\mathbf{e}^{+}$; representing the possible node representation of the target node in a link, with the corresponding source node representation being $\mathbf{z}_{i}$.
+We follow the same procedure for the hypervector encoding information about the edges that are absent, resulting in $\mathbf{Z}^{-}$.
 Precisely:
-$\mathbf{Z}^{+}$ $=\begin{bmatrix}\mathbf{z}_{1}\otimes\mathbf{e}^{+}\\ \vdots\\ \mathbf{z}_{N}\otimes\mathbf{e}^{+}\end{bmatrix}$ $\mathbf{Z}^{-}$ $=\begin{bmatrix}\mathbf{z}_{1}\otimes\mathbf{e}^{-}\\ \vdots\\ \mathbf{z}_{N}\otimes\mathbf{e}^{-}\end{bmatrix}$
+(3) $\begin{array}[]{cc}\mathbf{Z}^{+}=\begin{bmatrix}\mathbf{z}_{1}\otimes\mathbf{e}^{+}\\ \vdots\\ \mathbf{z}_{N}\otimes\mathbf{e}^{+}\end{bmatrix}&\mathbf{Z}^{-}=\begin{bmatrix}\mathbf{z}_{1}\otimes\mathbf{e}^{-}\\ \vdots\\ \mathbf{z}_{N}\otimes\mathbf{e}^{-}\end{bmatrix}\end{array}$
 Now, we find pairwise normalized distances between rows of $\mathbf{Z}^{+}$ and rows of $\mathbf{Z}$.
 This results in $\mathbf{D}^{+}=dist(\mathbf{Z}^{+},\mathbf{Z})\in[0,1]^{N\times N}$ , a matrix representing distances.
-If entry $(i,j)$ is close to zero, it indicates that the distance between row $i$ of $\mathbf{Z}^{+}$ is close to row $j$ in $\mathbf{Z}$, suggesting that $(i,j)$ might be a possible link.
-We perform the same process for $\mathbf{Z}^{-}$, resulting in the $\mathbf{D}^{-}=dist(\mathbf{Z}^{-},\mathbf{Z})\in[0,1]^{N\times N}$ matrix.
-However, an entry close to zero in $\mathbf{D}^{-}$ indicates a high possibility of not having an edge between the corresponding nodes.
-We now construct a prediction of the adjacency matrix represented by $\hat{\mathbf{A}}$, where entries $\hat{A}_{ij}\in[0,1]$.
-To construct the final reconstructed or predicted adjacency matrix of probabilities (based on distances), which factors in both the distances of having, or not having an edge using $\mathbf{Z}^{+}$ and $\mathbf{Z}^{-}$, we compute:
+If entry $(i,j)$ is close to zero, it indicates that the ”distance” between row $i$ of $\mathbf{Z}^{+}$ is close to row $j$ in $\mathbf{Z}$, suggesting the possibility of a link $i$ and $j$.
+Similarly, we find pairwise normalized distances between rows of $\mathbf{Z}^{-}$ and rows of $\mathbf{Z}$ yielding $\mathbf{D}^{-}=dist(\mathbf{Z}^{-},\mathbf{Z})$.
+An entry close to zero in $\mathbf{D}^{-}$ suggests the absence of an edge between the corresponding nodes.
+We now generate a prediction for the adjacency matrix, denoted by $\hat{\mathbf{A}}$, where $\hat{A}_{ij}\in[0,1]$.
+We predict values that fall within this interval because the evaluation metrics used are AUC-ROC and Average Precision.
+For real-world applications, an appropriate threshold can be selected depending on the specific use case to effectively interpret these predictions.
+To obtain the final predicted adjacency matrix of probabilities, which reflects the likelihood of an edge being present based on $\mathbf{D}^{+}$ and $\mathbf{D}^{-}$, we perform the following:
 $\hat{A}_{ij}=\begin{cases}\sigma\big{(}(1-D^{+}_{ij})+D^{-}_{ij}\big{)}&\text{if }D^{+}_{ij}<D^{-}_{ij}\\ \sigma\big{(}D^{+}_{ij}-(1-D^{-}_{ij})\big{)}&\text{else}\end{cases}$
@@
 We use the same logic when $D^{+}_{ij}\geq D^{-}_{ij}$ to obtain a value close to zero.
-5 Experiments
+5. Experiments
 We proceed to report results of extensive experiments comparing HDGL with a set of state-of-the-art node classification methods and link prediction methods, including several GNN models, in terms of both predictive performance and runtime.
-5.1 Data sets
-Since HDGL is designed in such a way that nodes with similar features and neighborhoods will have similar representations, we exclusively concentrate on homophilous graphs in this study, deferring the exploration of more complex cases involving heterophily graphs to future research.
-For our experiments, we utilize commonly used node labeling homophilous graph benchmarks: the CORA graph dataset [61], sourced from McCallum et al.
-[32]; the CiteSeer graph dataset [61], sourced from Giles et al.
-[16]; and the PubMed dataset [36].
+The code for HDGL can be found at: https://github.com/Abhishek-Dalvi410/HDGL.
+5.1. Data sets
+Since HDGL is designed to ensure that nodes with similar features and neighborhoods have comparable representations—often resulting in identical node labels—we focus exclusively on graphs with these characteristics as benchmark datasets for this study.
+For our experiments, we utilize commonly used node labeling graph benchmarks: the CORA graph dataset (Yang et al., 2016), sourced from McCallum et al.
+(2000); the CiteSeer graph dataset (Yang et al., 2016), sourced from Giles et al.
+(1998); and the PubMed dataset (Namata et al., 2012).
 Additionally, we include three more datasets from the Microsoft Academic Graph: Coauthor CS and Coauthor Physics from Shchur et al.
-[45], and the BlogCatalog Graph dataset from Huang et al.
-[20].
+(2018), and the BlogCatalog Graph dataset from Huang et al.
+(2017).
 Additionally, we introduce a new version of the DBLP dataset, built upon the works of Pan et al.
-[39].
+(2016).
 We remove nodes with no neighbors in the graph obtained from Pan et al.
-[39].
+(2016).
 The original node features from Pan et al.
-[39] consist of manuscript titles.
-We utilize BERT [11] to obtain embeddings for the titles and employ them as node features.
+(2016) consist of manuscript titles.
+We utilize BERT (Devlin et al., 2019) to obtain embeddings for the titles and employ them as node features.
 The data sets and their statistics are given in Table 1.
 We adhere to the Train/Validation/Test splits described in Yang et al.
-[61], Kipf and Welling [26], and Veličković et al.
-[53] for the CORA, CiteSeer, and PubMed datasets, respectively.
+(2016), Kipf and Welling (2017), and Veličković et al.
+(2018) for the CORA, CiteSeer, and PubMed datasets, respectively.
 For the Coauthor CS, Coauthor Physics, BlogCatalog, and DBLP datasets, we employ the data splitting strategy outlined in Shchur et al.
-[45].
-We test link prediction models on CORA, CiteSeer, and Pubmed datasets.
-These models are trained on incomplete data, retaining node features but omitting some edges.
-Validation and test sets include excluded edges and randomly selected unconnected node pairs.
-Evaluation assesses the models’ ability to classify both edges and non-edges.
-We follow splitting strategy from Kipf and Welling [25] work i.e Validation and test sets each contain 5% and 10% of citation links.
-5.2 Models
-For semi-supervised node classifcation task, we compare the performance of HDGL with that of several strong baselines Logistic Regression (LogReg), DeepWalk [42] and Label Propagation [64].
-We also directly compare HDGL with state-of-the-art models: Graph Convolutional Network (GCN) [26], Graph Attention Network (GAT) [53], and lastly, a fast GNN model, namely, Simplified Graph Convolutional Network (SGC) [57] which is essentially a logistic regression model trained using features extracted from $k$-hop neighborhoods of nodes as input.
-We also assess our model against RelHD [24], an existing HD algorithm used for node classification.
-For the transductive link prediction task, we evaluate our model against standard link prediction baselines:- DeepWalk [42] and Spectral clustering (SC) [48].
-We also compare our model against graph autoencoder models:- Graph Autoencoder (GAE) and Variational Graph Autoencoder (VGAE) from Kipf and Welling [25].
-5.3 HDGL Model Specifications
+(2018).
+We test transductive link prediction models on CORA, CiteSeer, and Pubmed datasets and with train/val/test splits from Kipf and Welling (2016) work i.e validation and test sets each contain 5% and 10% of links.
+Table 1. Dataset Summary.
+Dataset Nodes Edges Features Feature Type
+CORA 2708 5429 1433 $\{0,1\}$
+CiteSeer 3327 4732 3703 $\{0,1\}$
+PubMed 19717 44338 500 $(0,1)$
+BlogCatalog 5196 171743 8189 $\mathbb{Z}^{+}$
+Coauthor CS 18333 81894 6805 $\mathbb{Z}^{+}$
+Coauthor Physics 34493 247962 8415 $\mathbb{Z}^{+}$
+DBLP 17725 105781 768 $\mathbb{R}$
+5.2. Models
+For semi-supervised node classifcation task, we compare the performance of HDGL with that of several strong baselines Logistic Regression (LogReg), DeepWalk (Perozzi et al., 2014) and Label Propagation (Zhu and Ghahramani, 2002).
+We also directly compare HDGL with state-of-the-art models: Graph Convolutional Network (GCN) (Kipf and Welling, 2017), Graph Attention Network (GAT) (Veličković et al., 2018), and lastly, a fast GNN model, namely, Simplified Graph Convolutional Network (SGC) (Wu et al., 2019) which is essentially a logistic regression model trained using features extracted from $k$-hop neighborhoods of nodes as input.
+We also assess our model against RelHD (Kang et al., 2022), an existing HD algorithm used for node classification
+For the transductive link prediction task, we evaluate our model against standard link prediction baselines:- DeepWalk (Perozzi et al., 2014) and Spectral clustering (SC) (Tang and Liu, 2011).
+We also compare our model against graph autoencoder models:- Graph Autoencoder (GAE) and Variational Graph Autoencoder (VGAE) from Kipf and Welling (2016).
+5.3. HDGL Model Specifications
 To ensure that HDGL are directly comparable with most GNN architectures which typically use 2 layers, and hence aggregate information over 2-hop neighbors, we limit HDGL to aggregate information over at most 2-hops i.e HD node representations are calculated as per Eq.(2).
-Because Bundling of an even number of HD vectors can result in ties that would need to be broken using tie-breaking policy (See Section 3.2), we opt to randomly sample an odd number, specifically, 11 of the 1-hop neighbors and 21 of the 2-hop neighbors during the construction of HD node representations by HDGL for datasets except BlogCatalog.
+Because Bundling of an even number of HD vectors can result in ties that would need to be broken using tie-breaking policy (See Section 3.2), we opt to randomly sample odd number of neighbors, specifically, 11 of the 1-hop neighbors and 21 of the 2-hop neighbors during the construction of HD node representations by HDGL for datasets except BlogCatalog.
 For BlogCatalog we sample 21 of the 1-hop neighbors and 51 of the 2-hop neighbors since Blogcatalog graph is denser than the other datasets.
@@
 While these choices may seem arbitrary, it’s crucial to note that selecting dimensions between $20,000$ and $50,000$ typically suffices to maintain the desired properties of HD-computing.
-This assertion finds extensive support in the research from Kanerva [22, 23].
-Table 2: Average Test Set Accuracy and Standard Deviation (in Percentage) Over 10 Random Weight Initializations for Deep Learning Models and 10 Randomized Instantiations of RelHD and HDGL Across Datasets (* denotes datasets with $\{0,1\}$ features.
+This assertion finds extensive support in the research from Kanerva (1988, 2009).
+Table 2.
+Average Test Set Accuracy and Standard Deviation (in Percentage) Over 10 Random Weight Initializations for Deep Learning Models and 10 Randomized Instantiations of RelHD and HDGL Across Datasets (* denotes datasets with $\{0,1\}$ features.
 In these datasets, RelHD performs comparably to HDGL)
@@
 HDGL 79.5 $\pm$ 0.7 70.0 $\pm$ 0.6 76.8 $\pm$ 1.1 69.3 $\pm$ 0.5 88.9 $\pm$ 0.4 91.0 $\pm$ 0.9 68.7 $\pm$ 1.2
-5.4 Training and Evaluation
-Recall that iterative machine learning methods, including GNN typically make use of a validation data set that is distinct from the test data set for deciding when to terminate training, or for tuning the hyper-parameters.
-Hence, effectively, they make use of both training and validation data sets for building the predictive model and test data for evaluating the trained model’s performance.
-In contrast, learning methods such as HDGL which invole no iterative training do not need validation data.
-Hence, to ensure fair comparison between HDGL and other methods, we evaluate their performance on the training data, while letting the individual methods use training and validation data as appropriate for the algorithm in question.
-Thus, HDGL uses training and validation sets together for learning, whereas GNN use training data for training the model, validation data for early stopping and hyperparameter tuning.
+5.4. Training and Evaluation
+Iterative machine learning methods like GNN use separate validation data to stop training and tune hyperparameters, alongside training data.
+In contrast, methods such as HDGL which involve no iterative training do not need validation data.
+Hence, to ensure fair comparison, HDGL uses training and validation sets together for learning, whereas GNN use training data for training the model, validation data for early stopping and hyperparameter tuning.
 In the case of GNN methods that rely on iterative training, we choose the models with hyperparameters tuned to obtain optimal performance on validation data.
-For tuning the hyperparameters, we used grid search over a set of 12 configurations.
-In the case of GCN and GAT, the following hyperparameters were optimized the number of hidden units in the first layer, learning rate, and weight decay.
-In the case of GCN,networks with 32 and 64 hidden units were explored whereas in the case of GAT, 4 and 8 attention heads, each with 8 hidden units were explored.
-Both GCN and GAT were explored with weight decays in [1e-3, 1e-4, 5e-4].
-In the case of SGC which does not use any hidden layers, only the learning rate and weight decay parameters were tuned in [1e-3, 5e-4, 1e-4, 5e-5, 1e-5, 5e-6].
-GCN, GAT, and SGC were tuned for learning rates set to 1e-2 and 1e-3.
-The early stopping criterion optimized over choices of 25 and 50 epochs, respectively, for all three GNN models.
+We used grid search over 12 different configurations for tuning these hyperparameters.
+For GCN and GAT, we optimized the following hyperparameters: the number of hidden units in the first layer, learning rate, and weight decay.
+Specifically, GCN models were tested with 32 and 64 hidden units, while GAT models were tested with 4 and 8 attention heads, each having 8 hidden units.
+Both GCN and GAT models were evaluated with weight decays of 1e-3, 1e-4, and 5e-4.
+SGC, which does not have hidden layers, had its hyperparameters tuned for learning rate and weight decay.
+The weight decay tuning of SGC was over the following values 1e-3, 5e-4, 1e-4, 5e-5, 1e-5, and 5e-6.
+GCN, GAT, and SGC models were tuned with learning rates of 1e-2 and 1e-3, using early stopping criteria set to 25 epochs for the 1e-2 learning rate and 50 epochs for the 1e-3 learning rate.
 As LogReg, DeepWalk, and Label Prop represent well-established baselines, we do not conduct a hyperparameter search.
-For Label Propagation, we use 20 iterations of propagation (except for blogcatalog dataset where we use 10) and set the alpha factor  [^1] Alpha factor in Label Propagation determines the balance between existing and propagated labels.
+For Label Propagation, we use 20 iterations of propagation (except for blogcatalog dataset where we use 10) and set the alpha factor  [^1] Alpha factor in Label Propa determines the balance between existing and propagated labels.
 Higher alpha prioritizes existing labels; lower alpha favors propagated ones.
@@
 We sample such walks for all nodes in the graph and train DeepWalk model for 10 epochs and then employ logistic regression for classification, incorporating early stopping based on the validation set.
-Following Kang et al. [24], the RelHD algorithm was employed with 10,000 dimensional random basis HD vectors, as the performance of RelHD plateaus around this dimensionality.
+Following Kang et al. (2022), the RelHD algorithm was employed with 10,000 dimensional random basis HD vectors, as the performance of RelHD plateaus around this dimensionality.
 We report measured performance on the test data, averaged across 10 random selections of pertinent parameters.
@@
 For Cora, Citeseer, and Pubmed, we use predefined splits given in Yang et al.
-[61].
-For the link prediction task, we rely on the results directly obtained from the paper for GVAE from Kipf and Welling [25].
-5.5 Implementation Details
+(2016).
+For the link prediction task, we rely on the results directly obtained from the paper for GVAE from Kipf and Welling (2016).
+Figure 2.
+Comparison of Learning Times (in seconds) i.e Across Datasets for HDGL and RelHD with GCN, GAT and SGC with Hyperparameter Tuning with Exploration of Search Space with 12 configurations.
+Runtime for HDGL and RelHD exclusively includes only time Elapsed for learning label representations using train/validation data.
+[…]
+5.5. Implementation Details
 All experiments were conducted on Google Colab Pro with High-memory CPU.
-To ensure fair comparison, although not ideal for HDGL, we used the Python PyTorch framework [40], complemented by the Deep Graph Library (DGL) [54] for importing datasets and implementing deep learning models.
+To ensure fair comparison, although not ideal for HDGL, we used the PyTorch framework (Paszke et al., 2019), complemented by the Deep Graph Library (DGL) (Wang et al., 2019) for importing datasets and implementing deep learning models.
+Due to the lack of support for bitwise majority operations in PyTorch, we convert our vectors to the bipolar space and utilize signed addition for compatibility.
+As mentioned in Section 3.2, HD computing works in both binary and bipolar spaces, allowing us to switch between them as necessary.
 None of our experiments used parallel execution or GPUs.
-We have also developed a PyTorch version of RelHD, as the codebase was not publicly available and experiments detailed in Kang et al.
-[24] were reliant on hardware architecture implementations.
+We have also implemented PyTorch version of RelHD, as the experiments detailed in Kang et al.
+(2022) were reliant on hardware architecture implementations.
 We undertake this effort to compare with our model and use RelHD as a baseline in this study.
-5.6 Results
-Node Classification Performance.
-The results of our comparison of HDGL with the other methods are shown in Table 2.
-We observe that HDGL without the need for computationally expensive iterative training or hyperparameter optimization achieves node prediction accuracy that is competitive with those of state-of-the-art GNN models (GCN, GAT, SGC) across most of the benchmark data sets, in each case reaching performance that is within 1-2% of the best performing model.
-HDGL also consistently outperforming basic deep learning baseline methods (LogReg, DeepWalk, LabelProp).
-The notable advantage of HDGL over RelHD lies in its ability to accommodate features beyond the binary $\{0,1\}$ range, as evident from the results presented in Table 2, where HDGL significantly outperforms.
-This is particularly notable with datasets lacking $\{0,1\}$ features.
-However, when dealing with datasets featuring $\{0,1\}$ attributes like CORA and Citeseer, the performance of HDGL and RelHD appears nearly identical.
-To highlight the advantage of HDGL more prominently, we conducted experiments for node classification on the CORA and CiteSeer datasets where we reduced the node features to 100 dimensions using PCA [41], thereby trying to enforce continuous node features.
-As depicted in Figure 2, the impact is evident through a noticeable drop in performance for RelHD.
-Conversely, for other methods including HDGL, the performance remains relatively unchanged.
-Figure 2: Impact of performing PCA with a dimensionality of 100 on features for node classification accuracy.
-Figure 3: Comparison of Learning Times (in seconds) i.e Across Datasets for HDGL and RelHD with GCN, GAT and SGC with Hyperparameter Tuning with Exploration of Search Space with 12 configurations.
-Runtime for HDGL and RelHD exclusively includes only time Elapsed for learning label representations using train/validation data.
-Table 3: Link prediction results averaged over 10 Randomized Instantiations of HDGL and datasplits. Results for other models are from Kipf and Welling [25] paper.
-Methods Cora Citeseer PubMed
-AUC AP AUC AP AUC AP
-DeepWalk 0.831 0.850 0.805 0.836 0.844 0.841
-SC 0.846 0.885 0.791 0.826 0.849 0.888
-GAE 0.910 0.920 0.895 0.899 0.964 0.965
-VGAE 0.914 0.926 0.908 0.920 0.944 0.947
-HDGL 0.849 0.880 0.768 0.842 0.853 0.884
-Node Classification Run-times.
-Figure 3 compares the run-time for the node classification task between HDGL and RelHD (both of which requires no hyperparameter tuning) with those of GCN, GAT, and SGC (with hyperparameter tuning), alongside the run-time of HDGL.
-We observe that, the run-time of HDGL is substantially less than that of all the state-of-the-art GNNs.
-While in most cases, RelHD exhibits slightly or significantly worse run times compared to HDGL, it’s crucial to note that HDGL consistently delivers superior performance across various datasets when compared to RelHD.
-Table 4: Class Incremental Node Labeling Results.
+Table 3.
+Class Incremental Node Labeling Results.
 No hyperparameter grid search is used for GCN and SGC algorithms.
-Wall-time includes both training and inference of the model.
+Wall-time (in seconds) includes both training and inference of the model.
 HDGL does not require retraining from $t=2$.
@@
 Runtime 20.2s 0.3s 0.4s 0.4s 0.6s 38.3s 1.5s 1.7s 1.7s
-It is crucial to note that while we have used grid search on 12 potential setups for the deep learning model hyperparameter optimization of deep learning models typically explores a much larger space.
+5.6. Results
+Node Classification Performance.
+The results of our comparison of HDGL with the other methods are shown in Table 2.
+We observe that HDGL without the need for computationally expensive iterative training or hyperparameter optimization achieves node prediction accuracy that is competitive with those of state-of-the-art GNN models (GCN, GAT, SGC) across most of the benchmark data sets, in each case reaching performance that is within 1-2% of the best performing model.
+HDGL also consistently outperforming basic deep learning baseline methods (LogReg, DeepWalk, LabelProp).
+The notable advantage of HDGL over RelHD lies in its ability to accommodate features beyond the binary $\{0,1\}$ range, as evident from the results presented in Table 2, where HDGL significantly outperforms.
+However, when dealing with datasets featuring $\{0,1\}$ attributes like CORA and Citeseer, the performance of HDGL and RelHD appears nearly identical.
+Node Classification Run-times.
+Figure 2 compares the run-time for the node classification task between HDGL and RelHD (both of which requires no hyperparameter tuning) with those of GCN, GAT, and SGC (with hyperparameter tuning), alongside the run-time of HDGL.
+We observe that, the run-time of HDGL is substantially less than that of all the state-of-the-art GNNs.
+While in most cases, RelHD exhibits slightly or significantly worse run times compared to HDGL, it’s crucial to note that HDGL consistently delivers superior performance across various datasets when compared to RelHD.
+It is important to note that although we used grid search over 12 configurations for hyperparameter optimization, typically deep learning models explore a much larger parameter space.
 This expanded search naturally comes with increased runtime, especially when GPUs are utilized, further widening the energy efficiency gap between Graph Neural Networks (GNN) and traditional deep learning models like HDGL.
@@
 Nevertheless, beyond a certain threshold, the performance tends to plateau, as illustrated in Figure 3.
-Figure 4: Link Prediction Performance of HDGL under Various dimensionality configurations.
+Table 4. Link prediction results averaged over 10 Randomized Instantiations of HDGL and datasplits. Results for other models are from Kipf and Welling (2016) paper.
+Methods Cora Citeseer PubMed
+AUC AP AUC AP AUC AP
+DeepWalk 0.831 0.850 0.805 0.836 0.844 0.841
+SC 0.846 0.885 0.791 0.826 0.849 0.888
+GAE 0.910 0.920 0.895 0.899 0.964 0.965
+VGAE 0.914 0.926 0.908 0.920 0.944 0.947
+HDGL 0.849 0.880 0.768 0.842 0.853 0.884
+Figure 3. Link Prediction Performance of HDGL under Various Dimensionality Configurations.
+[…]
 HDGL in the Class-Incremental Learning Setting Consider a scenario with a dynamic graph where nodes and edges remain static, but the new node labels appear over time i.e initially, only a subset of nodes is labeled, belonging to two classes.
 Over time, new nodes receive labels; thus, introducing new classes.
-This is a node labeling instance of the class-incremental learning problem [52, 3, 51].
-In this setting, deep learning methods typically require either retraining on the data for previously learned classes along with the data for the new classes to avoid catastrophic forgetting of the previously learned classes [51].
+This is a node labeling instance of the class-incremental learning problem (Van de Ven et al., 2022; Belouadah et al., 2021; Tian et al., 2024).
+In this setting, deep learning methods typically require retraining on the data for previously learned classes along with the data for the new classes to avoid catastrophic forgetting of the previously learned classes (Tian et al., 2024).
 This retraining incurs significant additional runtime and computation.
@@
 We first split the nodes into train/val/test as in Shchur et al.
-[45].
-We mask the train/val/test such that at time $t=1$ only the labels $\{1,2\}$ are available; and at $t=2$ only the labels $\{1,2,3\}$ are only available; and so on such that at the final time step, all of the labels $\{1,\cdots,L\}$ are available.
+(2018).
+We mask the train/val/test such that at time $t=1$ only the labels $\{1,2\}$ are available; and at $t=2$ only the labels $\{1,2,3\}$ are available; and so on such that at the final time step, all labels $\{1,\cdots,L\}$ are available.
 To simplify matters for GNN models, we assume that the models are provided the list of $L$ possible labels although only a subset have been encountered in the training data.
@@
 In contrast, deep learning models like GCN and SGC mandate retraining for each timestep to accommodate new labels, resulting in comparatively slower runtimes.
-6 Conclusion
-We introduced HDGL, a novel and efficient algorithm for transductive learning on graphs using hyperdimensional representations.
-HDGL leverages the injectivity property of node representations of a family of Graph Neural Networks (GNNs) to map node features to the HD space and then uses HD operators such as bundling and binding to aggregate information from the local neighborhood of each node.
-The resulting latent node representations support both node classification and link prediction tasks, unlike typical deep learning methods, which often require separate models for these tasks.
-HDGL obviates the need for expensive iterative training of the sort required by deep learning methods.
-HDGL is also well-suited for class incremental learning where the model has to learn to effectively discriminate between a growing number of classes.
-Our experiments with widely used benchmark datasets show that on the node classification task, HDGL is competitive with the state-of-the-art GNN methods with respect to accuracy, at substantially reduced computational cost.
-Our experiments also show that the HD representation constructed by HDGL supports link prediction at accuracies comparable to that of DeepWalk and related methods.
-We conclude that HDGL offers a computationally efficient alternative to graph neural networks for node classification, especially in settings that call for class-incremental learning or in applications that demand high accuracy models at significantly lower computational cost and learning time than possible with the state-of-the-art graph neural network methods.
+6. Conclusion
+We introduced HDGL, a novel transductive learning algorithm for graphs using hyperdimensional representations.
+HDGL offers a computationally efficient alternative to graph neural networks by leveraging the injectivity property of node representations from Graph Neural Networks (GNNs).
+It uses HD operators such as bundling and binding to aggregate information from the local neighborhood of each node.
+The resulting latent node representations support both node classification and link prediction tasks.
+Furthermore, HDGL eliminates the need for iterative training, making it ideal for class-incremental learning and applications requiring high accuracy models at lower computational cost and learning time compared to traditional graph neural network methods.
+The advantages and efficacy of HDGL are validated by comprehensive experiments comparing it with state-of-the-art GNN methods.
+Acknowledgements.
+This work was funded in part by grants from the National Science Foundation (2226025), the National Center for Advancing Translational Sciences, and the National Institutes of Health (UL1 TR002014)